A Complete Bibliography of Publications in Mathematical Methods of Statistics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 April 2024 Version 1.30

Title word cross-reference

α [589, 628, 228]. AR(1) [108]. ARCH(p) [348]. β [226, 409, 663]. L [230]. D [281, 369, 294, 320, 148]. E [254]. F [643, 581]. G [331]. K [456, 674, 10, 668]. K_{n} [397]. L [66, 59, 582, 652, 656]. l^{n} [162]. L_{1} [344, 439, 466]. L_{2} [96]. L_{p} [49, 210, 285, 508]. M [127, 420, 364, 297, 286, 292]. L_{p} [564]. R [338, 639]. R^{+} [639]. R^{2} [247]. N [364, 140]. n: G [668]. P [536, 414, 294]. ϕ [345, 62]. ψ [661]. R [66, 59]. R^{d} [414]. s_{R}^{k−p} [95]. t [539]. τ [409]. U [461, 631, 589, 337]. V [337].

-balls [162]. -class [230]. -computation [331]. -convex [369]. -dependent [589].

-dimensional [320]. -Divergence [628, 345]. -divergences [581]. -estimation [466]. -estimators [127, 420, 297, 286, 292]. -losses [564]. -mean [210]. -method [344]. -minimax [96]. -Mixing [663, 409, 661, 226, 62, 228]. -nearest [397]. -norms [49]. -optimal [281, 254]. -out-of- [668]. -permutations [10]. -posterior [148, 294]. -Processes [631]. -Quantization [439]. -Record [674]. -regressive [226]. -sample [456, 140]. -Score [643]. -statistic [539, 589]. -statistics [461, 582, 337, 652, 656]. -tests [66, 59]. -value [294].

1 [506, 149]. 10 [246]. 11 [324]. 17 [434].

2 [24, 25, 513]. 2000c [196]. 2002e [246].

370 [434].

4 [111, 142].

60200a [111].

8 [196].

93m [149]. 95g [24, 25]. 96c [142]. 97c [111].

Abel [545]. absolute [125]. absolutely [29]. accuracy [364, 238, 87, 144, 108, 192]. accurate [561]. Acknowledgement [142]. Adaptation [481, 386, 64, 9]. adapting [112]. Adaptive [487, 498, 343, 214, 385, 431, 639, 237, 242, 493, 150, 187, 519, 409, 549, 268, 170, 644, 234, 155, 198, 473, 271, 320, 295, 199, 339, 231, 334, 203, 233, 119, 104, 222, 244, 305, 530, 564, 488, 509]. Addendum [142]. Additive [215]. Adjusted [626]. Admissibility [634]. Admissible [252]. affine [267]. against [455, 230]. aggregation [369, 370]. algebraically [189]. Algorithm [661, 558, 416]. Almost [325]. Alphabets [606]. Alternatives [629, 623, 515, 23, 17, 12, 24, 25, 92]. Analogue [671]. analogues [221, 246]. Analysis [677, 604, 156, 544, 299, 312, 494, 595]. analytic [102, 269, 300, 268, 195, 508]. Anderson [293]. anisotropic [457]. anisotropy [481]. ANOVA [245]. Application [472, 615, 291, 520, 287, 410, 403, 506, 513]. Applications [674, 665, 492, 536, 580, 438, 340, 412, 70, 139, 228, 273, 279, 337, 562]. Applying [647]. Approach [684, 614, 630, 596, 127, 349, 271, 461, 107, 372, 419, 270, 440, 263, 305, 255, 590, 27, 519, 14, 148]. Approximation [666, 624, 492, 536, 580, 559, 81, 381, 213, 48, 130, 29, 27, 42, 129, 512, 545, 416, 93, 146, 61, 579, 585, 541].

Approximations [438, 647, 456, 584]. arbitrary [7, 58]. ARCH [527, 406, 302]. Archimedean [265]. argument [587]. ARMA [406, 6]. ARMA-errors [6]. arrays [217, 228, 552]. Assessing [653, 657, 684, 599]. Associated [679, 166]. Association [347, 250]. assumptions [600, 219]. Asymptotic [626, 619, 613, 629, 515, 524, 555, 528, 617, 5, 388, 522, 48, 583, 72, 241, 26, 91, 471, 28, 296, 101, 10, 9, 19, 604, 350, 595, 243, 20, 585, 46, 73, 219, 403, 49, 177, 325, 211, 89, 261, 413, 179, 135, 483, 87, 144, 108, 18, 123, 286, 292, 355, 57, 138, 8, 121, 293, 106, 274, 181, 133, 34]. Asymptotically [437, 288, 164, 260, 102, 152, 23, 17, 12, 24, 25, 638, 298, 430, 275, 116, 123]. Asymptotics [503, 263, 68, 148, 194, 588]. Attraction [606]. Augmented [649]. auto [226]. autocorrelation [259]. Autoregression [618, 629, 623, 44, 76, 88, 115, 520, 87, 144, 350]. Autoregressive [609, 610, 36, 259, 528, 344, 109, 408, 19, 34, 441, 395]. average [138]. Averaging [42, 129, 27, 369, 572].

B [149]. Bahadur [127, 315, 179, 224, 92, 337, 293, 51]. balanced [544]. Balancing [608]. ballistic [522, 533]. balls [162]. Banach [439]. Bands [651]. Bandwidth [682]. Banks [497]. Baringhaus [92]. Barycenters [666]. Based [684, 669, 527, 366, 567, 1, 344, 263, 48, 476, 345, 249, 255, 557, 329, 494, 451, 565, 31]. bases [517]. Baxter [552]. Bayes [271, 72, 641, 79, 68, 99, 323, 106, 403]. Bayesian [596, 653, 657, 399, 586, 349, 445, 378, 281, 372, 419, 620, 323]. Behavior [619, 314, 89, 515, 91, 28, 585, 38, 403]. behaviour [325, 72, 135]. Bell [660]. Berkson [632]. Bernoulli [664]. Bernoullian [185]. best [512, 188]. betas [599]. between [497, 372, 419, 599, 94, 239, 645, 68, 4, 149]. BGWR [70]. bias [367, 504]. biased [52, 97]. Biclustering [612]. bilinear [28]. Binary [643, 609, 269, 185]. Binomial [620, 16]. Bivariate [605, 665, 255, 41]. blind [470]. blocks [266]. blockwise [231]. blurred [563]. Bochner [124]. bootstrap [48, 364, 266, 37, 32]. bootstrapped [492]. Bootstrapping [60, 539]. Borel [94]. Bound [676, 232, 583, 138]. Boundary [603, 224]. bounded [313]. Bounding [642]. Bounds [652, 656, 668, 399, 210, 538, 468, 475, 179, 579, 507, 32]. branching [70, 18]. Bridge [488]. brief [577]. Brownian [86, 111, 77, 675]. Burdick [174].

C. [435]. Calculate [647]. calibrated [32]. Can [664]. canonical [494, 595]. Cantelli [33]. Cardinal [379, 616]. case [503, 468, 475]. cases [72]. CCC [613]. CDF [287]. Censored [648, 637, 622, 444, 367, 255, 415, 20, 495, 553, 531]. Censoring [648, 184, 559, 343, 486, 489]. Central [631, 655, 70, 228, 341, 78]. cepstrum [483]. certain [570]. Cervonenkis [342]. chain [214, 336, 128, 191]. chains [299, 239]. Change [326, 356, 106, 151, 376, 288, 524, 438, 270, 482, 48, 346, 342, 132, 392, 576, 73]. Change-point [326, 106, 288, 438, 482, 48, 132, 392]. change-set [342]. Changes [630, 452]. channel [393, 392]. channels [403]. characteristic [366, 91, 257]. characteristics [347]. Characterization [39, 671, 201, 54, 550, 135, 464, 4, 149]. Characterizations [265, 646, 303, 361, 56, 160]. Characterizing [683, 587]. Chi [618, 5]. Chi-Square [618, 5]. choice [75, 117]. circle [254]. Circular [644, 668, 431, 83]. claims [380]. Class [673, 573, 420, 5, 116, 230, 120, 9]. Classes [597, 642, 320, 446, 342, 285]. classical [599]. Classification [643, 514, 172]. close [318]. clustering [439]. coefficient [376, 260, 233, 590, 216, 594]. Coefficients [610, 528]. Comment [434, 577, 435]. comments [118]. committed [477]. common [551]. compactly [358]. compare [561]. Comparing [317, 311]. comparison [90, 554]. complementary [303]. complexity [235]. component [430, 8]. components [487, 252, 571]. composite [236, 550]. comprehensive [554]. computation [224, 331]. computing [194]. concave [291]. Concentration [523, 531]. concept [294]. condition [356]. Conditional [631, 429, 107, 573, 517, 248, 39, 173]. Conditionally [609]. conditions [37, 355]. Confidence [651, 166, 85, 48, 284, 138, 427, 32]. Connected [659]. Consecutive [668]. Conservative [427]. Consistency [429, 667, 291, 207, 100, 110, 131, 395]. Consistent [360, 437, 158, 30, 183]. constancy [361]. constants [96]. constrained [525]. constraints [251, 535]. constructing [361]. construction [331]. Contacts [359, 365, 371, 377, 384, 391, 398, 405, 411, 417, 423, 428, 436, 443, 449, 454, 460, 467, 474, 480, 485, 491, 496, 500, 505, 511, 516, 521, 526, 532, 537, 542, 548]. contained [290]. Contaminated [619, 157]. context [533]. contiguous [515]. Continuous [670, 546, 211, 169, 555, 373, 518, 266, 120, 592]. continuous-time [518]. continuously [173]. contoured [504]. contrast [50]. Controlling [680, 272]. Convergence [309, 141, 71, 53, 272, 445, 199, 74, 35, 263, 208, 62, 576, 588, 105, 257, 579]. convex [360, 369, 197, 370]. Convolution [644, 301, 197, 8, 209]. convolutions [314]. Conway [683]. copula [387, 456, 492, 524, 580]. copulas [265, 602]. correct [32]. Correction [86, 111, 126, 442, 71]. corrections [440]. correctness [37]. correlated [247]. Correlation [638, 610, 574, 590, 442, 433]. correlations [4, 149]. corresponding [332]. corrupted [339]. Countable [606]. counterfactuals [331]. counting [158, 122]. coupling [3]. Covariance [434, 664, 435, 607, 569, 498, 56, 472, 459, 312, 560, 154, 404]. covariances [42]. covariate [495]. Covariates [677, 678, 634, 130]. covering [346]. Cox [158, 471, 216]. Cramér [246, 515, 582, 293, 221, 37]. Cramér-Rao [221]. Cramér-type [37]. Cramér-von [293]. criterion [497]. critical [95]. Crossing [603]. crossover [544]. cube [300]. cube-like [300]. cumulative [71, 53, 422, 495, 553]. Current [190, 291]. current-status [291]. Curve [650, 333, 309, 531, 601]. curves [203, 313]. CUSUM [73]. cut [529]. cut-off [529].

D [434, 654]. D-Optimal [654]. Darling [293]. Darmois [167]. Data [677, 678, 679, 626, 682, 608, 637, 622, 240, 663, 444, 43, 437, 1, 573, 291, 502, 256, 462, 322, 512, 545, 499, 415, 278, 375, 531, 52, 182]. Data-driven [240]. Data-Located [677]. decisions [566]. decomposable [67]. decomposition [465]. Deconvolution [632, 385, 598, 330, 564]. Deconvolving [358]. Deficiency [611, 66, 59, 229]. defined [247]. deformation [407]. degenerate [309, 337]. degrees [5]. delay [267]. delta [472, 410]. delta-method [472]. densities [487, 110, 549, 358]. Density [639, 621, 624, 628, 120, 604, 632, 77, 86, 111, 124, 338, 360, 282, 211, 295, 169, 555, 199, 517, 214, 385, 263, 316, 476, 113, 266, 457, 137, 161, 462, 409, 273, 330, 535, 426, 556, 370, 96, 234, 508, 484, 52]. Dependence [672, 665, 296, 255, 356, 38, 219]. Dependent [617, 645, 668, 166, 486, 385, 340, 589, 217, 408, 161, 322, 234, 151, 227, 223]. depending [173]. depth [503]. Derivatives [639, 161, 535]. Derived [670]. design [402, 309, 543, 563, 457, 6]. Designs [617, 654, 281, 254, 95]. detecting [270, 73]. Detection [633, 630, 452, 438, 162, 237, 242, 277, 321, 374, 425, 540, 534, 600, 151]. detectors [354]. determinant [134]. Deterministic [608]. Deviation [682, 624, 147, 210, 510, 499, 308, 507]. deviations [582, 67, 297, 132, 273, 389, 337, 355, 128, 227, 223]. difference [599, 159, 125]. differences [447]. different [317, 311]. differentiable [150]. differential [267]. diffusion [514, 260, 233, 289, 458, 336, 202, 9]. dimension [237, 242, 348]. Dimensional [675, 320, 259, 414, 336, 101, 209, 509]. dimensions [569, 137]. discounted [463]. Discrete [671, 453, 502, 236, 22, 13, 11, 575, 121]. Discriminant [299]. discrimination [239, 68]. Distance [63, 446, 476, 584, 30, 84, 451, 154, 302]. distances [581]. distinguished [264, 324]. distortions [172]. Distribution [618, 655, 617, 104, 630, 620, 314, 16, 45, 107, 115, 559, 486, 407, 333, 48, 584, 130, 80, 116, 153, 464, 396, 557, 18, 145, 576, 318, 565, 597, 8, 160, 594, 541, 38, 83, 525]. distribution-free [116]. Distributions [655, 670, 673, 645, 647, 671, 615, 646, 479, 107, 326, 163, 261, 453, 194, 72, 102, 39, 230, 126, 114, 551, 172, 21, 123, 318, 58, 4, 504, 196, 171, 197, 547, 201, 206, 252, 191, 149, 133]. Divergence [628, 645, 283, 345]. divergences [581]. divisible [126, 114]. does [125]. Domains [606, 300]. Donsker [446]. Donsker-classes [446]. Double [659]. Doubly [673]. down [576]. drift [406, 233, 289, 458]. Driven [675, 256, 240]. Dutilleul [435]. Dvoretzky [651]. dyadic [481]. dynamical [225, 165].

edges [269]. Edgeworth [381, 335, 364].

Effects [649, 103, 549]. Efficiency [565, 182, 315, 357, 522, 583, 179, 193, 67, 224, 58, 92, 337, 355, 274, 134]. Efficient [300, 353, 451, 191, 258, 594, 441, 164, 275, 260, 102, 40]. Efron [37]. ellipsoids [65, 427]. Elliptical [655, 673]. elliptically [504]. EM-estimator [41]. emperors [577]. Empirical [378, 682, 618, 631, 98, 396, 627, 99, 642, 276, 314, 399, 49, 271, 115, 200, 262, 456, 492, 580, 424, 363, 217, 520, 335, 341, 280, 249, 28, 506, 513, 329, 287, 250, 207, 257, 71, 53, 38]. Entropic [606]. entropy [397]. environment [522, 533, 452]. equality [472, 116]. equals [2]. Equation [647, 437]. Equations [675, 275, 267, 159]. equivalence [241]. ergodic [555, 573, 260, 447, 307]. ergodicity [400]. Errata [24, 25, 22]. Erratum [246, 34, 196, 324]. Error [677, 637, 470, 396, 317, 311, 58, 8, 93, 146, 207, 594]. Errors [617, 676, 661, 614, 632, 226, 477, 306, 310, 426, 6]. Errors-in-Variables [614, 226]. estimability [330]. Estimate [683, 100, 8]. estimates [1, 36, 115, 288, 573, 482, 72, 30, 29, 3, 137, 302, 143, 97]. Estimating [679, 608, 458, 535, 437, 42]. Estimation [527, 272, 613, 211, 259, 283, 407, 639, 470, 113, 7, 543, 621, 374, 425, 551, 159, 609, 204, 307, 336, 628, 434, 426, 615, 463, 404, 435, 665, 663, 525, 607, 395, 349, 338, 360, 166, 178, 320, 282, 498, 295, 200, 169, 343, 372, 419, 199, 486, 339, 164, 231, 517, 420, 490, 157, 275, 214, 226, 431, 373, 260, 357, 247, 333, 453, 502, 203, 533, 518, 446, 158, 408, 476, 309, 233, 289, 583, 119, 102, 104, 80, 205, 222, 305, 457, 367, 354, 220, 118, 321, 47, 493, 284, 459, 82]. estimation [285, 530, 161, 590, 165, 202, 120, 123, 462, 150, 187, 192, 519, 409, 300, 549, 19, 34, 268, 588, 301, 353, 375, 287, 380, 84, 350, 494, 451, 105, 170, 99, 121, 216, 356, 556, 106, 40, 191, 258, 594, 96, 441, 541, 348, 234, 508, 198, 276, 466, 183, 235, 601]. Estimator [604, 429, 430, 559, 332, 413, 522, 26, 477, 79, 483, 131, 101, 229, 499, 572, 243, 20, 41]. Estimators [672, 667, 617, 681, 251, 133, 124, 127, 382, 325, 211, 88, 524, 555, 420, 263, 316, 48, 90, 476, 98, 570, 266, 135, 297, 112, 238, 396, 87, 144, 108, 317, 311, 286, 292, 273, 415, 442, 433, 397, 389, 58, 597, 370, 484, 488, 227, 52, 403, 50]. evaluated [38]. Evaluation [679, 238]. event [278]. event-history [278]. events [21]. Exact [339, 647, 194]. Examples [419]. excursions [75, 117]. Existence [683, 400, 208]. expansion [261, 363, 179, 335, 364, 471]. expansions [388, 381, 26, 19, 34]. Expectation [464, 642]. Expectations [652, 656, 173]. expected [463]. experiments [103, 290, 128, 6]. Explicit [544, 557]. explosive [144]. exponent [133]. Exponential [538, 603, 303, 418, 593, 54, 523, 574, 551, 464, 189, 4, 565, 149, 525, 571]. exponentiality [577, 230, 355, 565]. exponentiated [353]. expressed [452]. expressions [557]. Extended [560, 601]. extension [167]. extensions [574]. Extremal [566, 16]. Extreme [672, 272, 559]. extremum [304].

factor [390]. factorial [95]. factorization [586]. Families [603, 303, 418, 593, 54, 2, 189, 68]. family [30, 442, 433, 547]. feedback [129]. field [15, 46]. fields [429, 247, 78, 75, 117, 279]. filtered [290]. filtering [65, 271, 152]. filters [575]. finite [518, 386, 47, 299, 128]. First [647, 408]. Fisher [318, 584, 63]. Fisher-type [63]. Fit [629, 412, 249, 193, 421, 568, 145, 240]. fitting [327]. fixed [166, 87, 144]. forms [507, 562]. formula [66, 59, 331]. foundations [599]. Fourier [254, 363]. fractional [468, 475, 95]. framework [540]. Fredholm [675]. free [116]. freedom [5]. Function [618, 667, 674, 670, 654, 669, 662, 314, 366, 1, 115, 339, 104, 457, 354, 91, 530, 101, 322, 499, 415, 432, 572, 463, 541, 495, 553, 41]. Functional [682, 212, 676, 168, 105, 319, 573, 431, 35, 587, 158, 341, 493, 202, 499]. functionals [77, 86, 111, 49, 261, 113, 305, 26, 477, 28, 186, 493, 204, 123, 501, 506, 513, 576, 301, 198]. Functions [624, 675, 567, 210, 54, 80, 321, 374, 425, 307, 568, 150, 187, 501, 506, 513, 300, 545, 268, 504, 170, 197, 195, 257, 71, 53, 40, 96, 308, 508, 38].

gain [146]. Gamma [648, 382]. GARCH [400, 613, 276]. Gaussian [349, 536, 559, 528, 666, 232, 470, 305, 393, 483, 75, 117, 27, 129, 501, 15, 195, 160, 234, 73]. General [667, 603, 534, 127, 286, 292, 318, 206, 46, 395, 501]. generalization [56]. generalizations [293]. Generalized [605, 382, 413, 193, 369, 486, 130, 447, 79, 464, 557, 154, 160, 308]. generated [303]. Generating [674, 680, 662, 432]. Geometric [669, 400]. Gini [599]. given [581, 380]. Glivenko [33]. global [57]. GM [520]. GM-testing [520]. GMANOVA [560]. Good [659]. Goodness [145, 412, 249, 193, 421, 568]. Goodness-of-fit [145, 412, 249, 193, 421, 568]. Grouped [401]. growing [237, 242]. growth [601]. guarantee [551].

H [435]. half [497]. hazard [602, 1, 255, 554]. hazard-based [255]. Heavy [625, 615, 559, 279]. heavy-tailed [559, 279]. height [153]. Hellinger [30]. Help [359, 365, 371, 377, 384, 391, 398, 405, 411, 417, 423, 428, 436, 443, 449, 454, 460, 467, 474, 480, 485, 491, 496, 500, 505, 511, 516, 521, 526, 532, 537, 542, 548]. Henze [92]. heterogeneity [278]. heteroscedastic [200, 262, 509]. heteroscedasticity [488]. hidden [152, 336, 403]. High [75, 117, 374, 336, 509, 219]. high-dimensional [336, 509]. high-variable [374]. Higher [177, 561, 137]. Hilbert [340]. Hilbert-valued [340]. Hilbertian [250]. Hill [413, 587]. Histogram [444]. history [278]. Hölder [257, 320, 642]. Hölderian [341]. homodyne [354, 375]. homogeneity [15]. hybrid [438]. Hyperbolic [670]. hypotheses [43, 406, 236, 550, 225, 145, 195, 566, 148]. Hypothesis [630, 319, 200, 328, 23, 17, 12, 24, 25, 126, 114, 306, 310, 368, 155]. hypothetical [172].

ideal [354]. identifiability [330]. identities [464]. identity [569, 56]. II [24, 66, 475, 419, 103, 17, 242, 310, 22, 13, 117, 144, 317, 506, 513, 450, 292, 227]. Iid [652, 656]. III [23]. Ill [676, 529, 543]. ill-posed [529, 543]. Ill-Posedness [676]. Image [393, 269, 204]. images [185]. immigration [18]. Implication [676]. imply [125]. improve [192]. Improved [672, 597]. incidental [143]. incomplete [1, 326]. inconsistent [175]. Increase [608]. increment [581]. Independence [125, 387, 469, 163, 201, 274, 383, 298, 473]. independent [125, 189, 179]. Index [672, 608, 497, 559, 519]. Indexed [642]. indirect [570, 82, 451]. individual [278]. induced [212]. Inequalities [340, 327, 351, 523, 562]. Inequality [659, 650, 586, 221, 246, 552]. Inference [660, 678, 612, 345, 424, 518, 557, 451]. Inferential [650]. infinite [115, 259, 101, 209]. infinite-dimensional [259, 101, 209]. infinitely [126, 114, 150]. Inflated [660, 678]. influence [544, 313]. Information [674, 270, 662, 326, 318, 290, 63]. Information-theoretic [270]. inhomogeneity [481]. inhomogeneous [458, 64]. input [275]. inputs [385]. insurance [463]. Integral [675, 194]. integrated [580, 412, 249, 224, 8, 61, 198]. integration [215]. intensity [567, 538, 368, 122]. intermediate [479, 193, 67]. Interpolation [616, 512, 578]. Interval [184, 268]. intervals [166, 48, 138]. Intraclass [664, 442, 433]. Introduced [611]. invariance [250, 390, 151]. Invariant [634, 191, 143, 140]. Inverse [633, 626, 627, 540, 352, 534, 287, 160]. inversion [14]. Inverted [655]. invertible [258]. invitation [220]. Involving [682, 251]. ion [403]. isomorphism [94]. isotropic [584]. iterated [168]. Iteration [670].

Jacobi [545]. Jensen [659]. Jittering [604]. Johann [176]. joint [77, 86, 111, 602, 153]. jump [336]. jumps [538, 170, 394].

Kaplan [248, 33, 229]. Kernel [1, 667, 373, 640, 604, 97, 124, 282, 211, 215, 316, 137, 389, 61, 227, 52]. Kernel-Type [667]. Kernels [667, 154]. Khmaladze [412]. Kiefer [651, 127, 168, 51]. known [430]. Kolmogorov [596]. Kronecker [434, 435, 459, 404]. Kullback [283].

LAD [20]. Laguerre [598]. LAN [324, 179, 264]. LAN-representations [324, 264]. Laplace [263, 194, 329]. Large [349, 147, 67, 297, 624, 273, 499, 389, 355, 308, 569, 245, 326, 5, 510, 21, 132, 337, 128, 410]. Lasso [401, 488, 509]. lattice [163]. law [573, 587, 168, 410]. laws [303, 497, 415, 329, 353]. leading [126, 114]. Least [617, 235, 527, 26, 131, 307, 108]. least-squares [108]. Lee [435]. left [322]. Leibler [283]. lemma [124]. length [153, 138, 52]. lengths [577]. level [641, 75, 117, 64]. Lévy [303, 424, 463]. life [230]. Lifetime [645]. Lifetimes [658, 367]. like [300]. Likelihood [607, 315, 166, 524, 35, 522, 533, 476, 471, 135, 47, 396, 290, 571, 97]. Lilliefors [355].

Limit [605, 631, 279, 479, 468, 475, 35, 212, 70, 333, 341, 280, 78, 228, 415, 196, 171]. Limiting [45, 573, 38]. limits [136]. line [441]. Linear [617, 614, 654, 370, 465, 529, 431, 446, 130, 205, 244, 305, 543, 563, 280, 471, 159, 493, 110, 82, 37, 462, 549, 301, 390, 595, 252, 356, 106, 427, 258, 240]. Linearity [2]. linearly [560]. Local [319, 612, 682, 455, 629, 623, 169, 123, 64, 651, 503, 468, 475, 346, 122, 92, 355, 57, 146, 356, 293]. Locally [103, 217, 243]. Located [677]. Location [683, 673, 175, 2, 68, 525, 466]. Location-Scale [673, 68]. logarithm [415, 168]. Logistic [680, 345, 208]. long [280, 296]. long-range [296]. Longitudinal [626, 437]. look [539]. Loss [628, 197]. losses [222, 564]. Lower [85, 676, 210, 583, 464, 557, 179].

M [434]. MA [518]. Manifolds [667, 372, 419]. Mann [589]. many [130, 321]. MAP [641]. maps [372, 419]. marginal [549]. Marginalized [678]. marginals [116]. Markov [157, 214, 152, 98, 299, 336, 239, 636, 128, 191, 395, 403]. Markovian [461]. martingale [447, 146]. martingales [70, 223]. Math [111, 246, 24, 25, 142, 196, 324, 149]. Mathematical [434]. Matrix [655, 681, 607, 569, 586, 598, 39, 318, 560]. Maximal [665]. maximin [281, 550]. Maximum [533, 47, 607, 166, 524, 522, 476, 135, 97]. Maxiset [338, 633, 349]. maxisets [323]. Maxwell [683]. Mean [619, 55, 142, 361, 418, 593, 210, 334, 304, 328, 335, 364, 29, 2, 79, 18, 576, 32]. mean-value [29]. Means [634, 81, 55, 142]. Measure [645]. Measurement [677, 676, 221, 246, 317, 311]. measures [386, 345, 141, 69]. median [36, 55, 142]. medians [55, 142]. Meier [248, 33, 229]. memory [280]. Method [612, 344, 215, 231, 263, 194, 472, 523, 3, 459, 84, 541, 410, 50].

Methods [111, 361, 246, 24, 25, 434, 142, 196, 324, 149, 578]. minimal [45]. Minimax [236, 614, 457, 354, 162, 421, 478, 132, 448, 450, 600, 301, 644, 140, 52, 65, 80, 205, 23, 17, 12, 24, 25, 123, 150, 350, 96, 298, 473]. minimaxity [87, 144]. minimizers [399]. Minimum [446, 30, 302, 476, 581, 84, 451, 50]. mirror [369]. Miscategorized [677]. Mises [584, 515, 293]. Mises-type [515]. Missing [678, 626, 182]. Mitscherlich [654]. Mixed [675, 232, 549, 588]. mixed-effects [549]. mixed-rates [588]. Mixing [631, 661, 663, 226, 409, 62, 228, 556]. Mixture [673, 487, 430, 590]. ML [683]. MLE [528, 175, 208]. Mode [453, 622, 573, 471]. Model [660, 683, 612, 620, 649, 60, 445, 498, 430, 344, 270, 157, 226, 158, 289, 152, 305, 393, 459, 110, 101, 317, 311, 519, 122, 301, 442, 433, 560, 426, 216, 427, 594, 302, 348, 465, 276, 395]. modeling [323]. Models [678, 613, 655, 617, 614, 434, 636, 680, 404, 435, 665, 646, 382, 36, 44, 406, 387, 524, 529, 431, 254, 261, 130, 205, 244, 269, 543, 563, 185, 345, 267, 544, 255, 493, 3, 204, 396, 590, 37, 188, 549, 278, 251, 390, 20, 40, 191, 394, 240, 546, 509, 466, 97, 403, 182, 601]. Moderate [682, 227, 223, 582, 389]. Modification [677, 416]. Modified [41, 124]. modulated [336]. Moment [588, 353, 432, 287, 547]. moment-empirical [287]. moment-type [353]. Moments [673, 606, 400, 468, 475, 518]. Monotone [402, 547, 231]. Monotonicity [684, 641, 362]. Monro [388, 661]. most [103]. Motion [675, 77, 86, 111]. moving [579]. MR1202799 [149]. MR1257978 [25]. MR1257983 [24]. MR1324693 [142]. MR1372016 [111]. MR1692715 [196]. MR1841806 [246]. MR1979746 [324]. Multi [641, 393, 392, 465]. multi-channel [393, 392]. Multi-level [641]. multi-way [465]. Multichannel [277]. multidimensional [536, 540]. Multilevel [677]. multinomial [67]. multiparameter [76, 88, 72]. Multiple [630, 524, 590, 588, 595]. multiple-set [595]. multiplicative [122, 416]. Multiplicity [647]. multiscale [333]. Multivariate [653, 657, 613, 555, 332, 673, 624, 459, 274, 366, 43, 178, 382, 469, 515, 438, 203, 584, 116, 421, 79, 110, 172, 530, 161, 187, 300, 389, 572, 15, 257, 383].

N [459]. natural [303, 418, 593, 54]. nearest [397]. nearly [108]. Necessary [653, 657, 148]. NEF [361]. Negative [620]. neighbor [397]. Nevzorov [149]. Neyman [256]. no [111, 246, 24, 25, 142, 196, 324]. noise [528, 339, 232, 7, 185, 393, 225, 165, 202, 600, 195]. noises [408]. noisy [375]. Nominal [677]. Non [468, 475, 655, 654, 330, 636, 208, 224, 337]. Non-Central [655]. non-degenerate [337]. Non-estimability [330]. non-existence [208]. Non-Linear [654]. non-standard [224]. Non-uniform [468, 475]. nonasymptotic [184]. noncentral [280]. noncentrality [590]. Noncommutative [221, 246]. nonhomogeneous [567, 73]. Nonlinear [609, 344, 490, 261, 26, 296, 317, 311, 84, 494]. nonnegative [586, 328]. Nonparametric [653, 657, 376, 643, 502, 476, 304, 289, 367, 640, 306, 310, 368, 186, 284, 161, 202, 578, 375, 380, 216, 556, 429, 60, 596, 349, 43, 362, 44, 76, 88, 343, 424, 74, 379, 263, 333, 402, 328, 233, 119, 241, 23, 17, 12, 24, 25, 277, 421, 478, 238, 172, 396, 139, 187, 228, 499, 415, 350, 312, 195, 155, 97, 32, 182]. nonregular [72]. norm [320, 222]. Normal [130, 29, 107, 430, 270, 79, 123, 318, 68, 561, 133, 592]. Normality [653, 657, 679, 366, 49, 413, 522, 39, 125, 101, 432, 57, 20, 46]. normalization [479]. normalized [218, 253]. normed [446]. norms [49]. note [281, 312, 484, 331]. NPMLE [182]. nuisance [177, 543, 188, 143]. null [236]. number [245, 5, 21, 18]. numbers [346, 189, 410]. Numerical [675].

observations [179, 82, 307, 161]. Occupation [636]. off [529]. offspring [18]. Oja [635]. Omnibus [329]. on-line [441]. One [414, 497, 66, 59, 304, 328, 2, 312, 313]. One-dimensional [414]. one-half [497]. one-sample [66]. one-sided [304, 328, 312, 313]. online [572]. Operator [567, 259, 598, 352]. Operator-based [567]. operators [472]. Optimal [210, 74, 275, 643, 639, 341, 654, 285, 578, 616, 122, 638, 96, 6, 437, 281, 288, 254, 152, 570, 575, 75, 117, 139, 545, 239]. optimality [177, 5, 9, 92, 355, 196, 171, 243, 293]. Optimally [313]. Oracle [445, 586, 327, 351, 305, 519]. oracles [231]. Order [605, 684, 635, 669, 602, 479, 177, 357, 212, 520, 408, 583, 80, 205, 266, 574, 464, 37, 239, 561, 554, 219, 32]. Ordered [646, 551]. ordering [571]. ordinal [465]. origin [360]. orthogonal [273]. Ostensibly [625]. Outliers [618, 623, 625, 455].

P [435]. PAC [399]. paper [149]. parallel [554, 571]. Parameter [82, 272, 166, 558, 157, 85, 522, 408, 26, 118, 458, 2, 75, 117, 101, 19, 34, 99, 121, 209, 348]. Parameters [609, 177, 406, 270, 587, 7, 173, 551, 159, 87, 144, 590, 188, 251, 597, 427, 441, 143, 525, 546, 276, 466]. Parametric [420, 622, 524, 533, 267, 145, 278, 148]. parametrization [593]. Pareto [597]. Part [506, 513]. partial [340, 254, 205, 244, 576]. Past [658]. Path [486]. Path-dependent [486]. paths [295, 514]. PDE [135]. PDE’s [121]. Pearson [618, 623, 638]. Penalization [612]. Penalized [666, 231]. penalties [235]. penalty [256, 463]. period [339]. Periodic [613, 400, 339, 269]. periodically [247]. periodicity [458]. periods [3]. permutation [390, 181]. permutations [10]. perspectives [512]. Perturbation [143]. perturbed [38]. Pfanzagl [176]. Phillips [563]. Piecewise [462, 481]. Pinsker [232]. plots [536, 414]. plug [483]. plug-in [483]. Point [633, 184, 288, 326, 438, 482, 48, 30, 414, 132, 392, 145, 356, 106, 38, 151]. points [503, 524, 336]. Pointwise [530, 199, 309, 457, 535, 243, 426]. Poisson [678, 567, 683, 583, 368]. polynomial [481, 418, 281, 327, 21]. polynomials [156, 556]. polytomous [60, 345]. population [47, 188]. Populations [652, 656, 116, 561]. posed [529, 543]. Posedness [676]. positive [501, 506, 513, 329]. possibly [444, 115]. Post [643]. Post-Processing [643]. posterior [445, 72, 130, 2, 148, 294]. Power [629, 623, 479, 31]. powerful [103]. pp [434]. precedence [414]. Precision [681]. predictable [558]. predicting [110]. Prediction [344, 637, 259, 470]. Predictive [620, 628]. Preliminary [442, 433]. presence [343, 415]. Principle [682, 147, 308]. principles [389, 250, 151]. prior [445]. priority [142]. priors [349]. Probabilities [603, 636, 380]. Probability [626, 647, 208, 201, 96]. problem [288, 326, 456, 164, 316, 482, 478, 346, 342, 392, 139, 174, 206, 95, 140]. Problems [633, 361, 177, 304, 126, 114, 540, 132, 224, 352, 534, 287, 121, 14, 566]. procedures [44, 570, 345, 561, 579, 585, 73]. Process [627, 610, 567, 528, 438, 388, 248, 502, 213, 587, 408, 280, 249, 28, 368, 87, 144, 18, 122, 250, 146, 234, 73, 422, 489]. Processes [682, 631, 13, 11, 642, 527, 147, 400, 211, 200, 262, 259, 456, 492, 536, 555, 580, 424, 490, 35, 373, 412, 70, 81, 109, 217, 520, 518, 158, 233, 289, 583, 98, 266, 341, 118, 414, 108, 568, 202, 120, 506, 513, 409, 9, 145, 576, 168, 258, 276, 22]. Processing [643]. Product [434, 435, 425, 94, 459, 141, 69, 404]. profile [471]. projection [445]. Properties [613, 524, 528, 266, 483, 9, 287, 105, 547, 31, 73, 219]. property [16, 565]. proportional [265]. pseudo [524]. Purely [144]. purity [375].

QML [613]. quadratic [261, 26, 507, 151, 562, 105]. Quantile [669, 663, 429, 341, 322, 71, 53, 308, 422, 495, 553, 489]. Quantiles [672, 89, 61]. quantitative [410]. Quantization [439, 207, 206]. quantizations [252]. quantum [354, 220, 221, 246, 375]. Quasi [197, 586, 465]. quasi-Bayesian [586]. Quasi-convex [197]. quasi-symmetry [465]. questions [570].

Radon [478]. Random [659, 605, 608, 661, 664, 610, 649, 646, 429, 156, 210, 559, 407, 103, 247, 402, 522, 533, 472, 584, 452, 78, 125, 501, 506, 513, 512, 62, 549, 55, 142, 279, 380, 591, 390, 146, 46, 38, 489]. random-effects [549]. randomized [399]. range [296]. rank [147, 115, 103, 91, 274, 601]. ranking [312]. Rao [246, 221]. rare [21]. Rate [35, 208, 579, 445, 602, 263, 122, 554]. Rates [667, 643, 199, 74, 309, 141, 588, 128, 96]. Ratio [413, 615, 315, 35, 547, 571]. ratios [290]. Rayleigh [557]. Real [663]. Real-Valued [663]. reconstruction [316, 393]. Record [674, 662, 557]. records [4, 149]. recovery [269, 285, 139]. Rectangular [671]. Recursive [558]. reduced [601]. regenerative [3]. regions [85]. Regression [660, 678, 418, 667, 648, 676, 640, 654, 622, 627, 578, 680, 60, 303, 361, 281, 362, 555, 343, 74, 379, 373, 254, 261, 291, 402, 158, 236, 7, 119, 241, 26, 457, 345, 100, 208, 421, 296, 396, 131, 307, 101, 37, 139, 499, 415, 389, 572, 84, 190, 426, 195, 160, 20, 71, 53, 40, 594, 6, 422, 495, 553]. regressive [226]. regret [140]. Regular [617, 214, 29, 206, 133]. regularity [30].

Regularization [352, 534]. regularizations [529, 563]. regularized [588]. reign [577]. related [361, 56, 327, 351, 164, 482, 304, 478, 31]. relation [224]. Relative [637, 229]. Reliability [668]. remark [183]. Renewal [109, 461, 502]. representation [248, 322, 51]. Representations [591, 127, 264, 324]. Resampling [611]. Residual [520, 645, 262]. Residuals [618, 627, 115, 312, 160]. resolution [64]. respect [229, 160]. Response [677, 435, 553]. Responses [626]. Restoration [185]. restricted [85]. result [554]. Results [650, 555, 573, 401, 109, 217, 273, 499, 31]. reversed [602]. Révész [572]. review [593]. revisited [593, 448, 450]. reward [502]. Riemannian [372, 419]. right [343]. Risk [570, 399, 316, 132, 68, 504, 106, 128]. Risk-optimal [570]. risks [463]. Robbins [388, 661]. Robust [136, 157, 84, 79, 276]. Robustness [172, 455, 575]. Rosen [434]. Roy [435, 459]. ruin [380]. rules [504, 323]. RWRE [621].

S [434]. S. [459]. Saddlepoint [584]. Sample [679, 605, 619, 295, 596, 361, 418, 66, 59, 89, 456, 2, 94, 380, 337, 140, 148, 592]. sampled [462]. Samples [680, 326, 15, 383]. sampling [123, 121, 134, 97]. saturated [95]. Scale [673, 652, 656, 551, 68, 99, 40]. scheme [236, 67]. schemes [22, 13, 11, 27, 129, 534]. Score [643, 568, 31]. score-functions [568]. seasonal [280]. Second [205, 266, 239, 357, 518, 583, 80, 37]. Second-order [266, 357]. Selecting [649]. Selection [640, 188, 444, 481, 445, 498, 289, 367, 426]. Self [253, 218]. Self-normalized [253, 218]. Semi [622, 524, 98]. semi-Markov [98]. Semi-Parametric [622, 524]. semimartingale [93, 394].

Semiparametric [430, 226, 225, 165, 357, 191, 394]. Separation [680]. Sequences [631, 340, 589, 477, 280, 243, 323, 151]. Sequential [233, 269, 147, 262, 408, 289, 19, 34, 239]. serial [91]. Serials [672]. Series [609, 558, 200, 262, 215, 470, 100, 483, 273, 105]. set [284, 346, 342, 595, 566]. set-valued [566]. setting [501]. setup [166]. shape [339, 137]. Sharp [320, 676, 507, 231, 203]. Sheppard [440]. Shift [390, 357, 92]. shifts [407]. Shiryaev [73]. shorth [325, 153]. shortt [213]. shrinkage [357]. sided [304, 328, 312, 313]. sieves [50]. sign [44, 76, 88, 455]. Signal [633, 540, 162, 237, 242, 277, 82, 534, 600]. signed [386]. simple [208, 91]. simplicial [503]. simulation [476, 451]. simulation-based [476, 451]. simultaneous [90]. Sine [670]. single [519]. single-index [519]. Size [605, 2, 416, 148]. sizes [442, 433]. skew [432]. Skewness [679, 635, 447]. Skitovich [167]. slightly [452]. slippage [140]. Sloan [374]. Small [632, 587, 225, 306, 310, 165, 202]. Smirnov [596]. smooth [288, 321, 204, 307, 101, 187]. smoothed [229]. smoothing [351, 104, 244, 75, 117, 122, 356, 97]. smoothness [378, 112, 64]. Sobolev [285]. Solution [647, 675, 139, 140]. solutions [437, 267]. Some [469, 580, 401, 217, 126, 114, 512, 287, 646, 387, 275, 70, 29, 574, 28, 110, 202, 224, 273, 504, 160, 191, 151, 227, 223, 601, 227]. Space [675, 107, 386, 374, 94]. spaces [85, 446, 425, 439, 209]. spacing [45]. spacings [10, 219]. sparse [323, 509]. Spatial [282, 663, 373, 347]. spatially [119, 155]. special [565]. specification [394]. Spectral [529, 247, 518, 234]. speed [272]. sphere [15]. sphericity [569]. spite [330]. spline [131]. splines [351, 379]. Square [618, 5]. squared [8]. Squares [617, 527, 26, 131, 307, 108, 235]. Srivastava [434]. stability [497]. stable [303, 497, 118, 318, 329, 87]. standard [224]. State [636, 375]. Stationarity [400, 527]. Stationary [631, 503, 555, 573, 528, 267, 78, 447, 299, 307, 409]. Statist [111, 246, 24, 25, 142, 196, 324, 149]. Statistic [638, 539, 589, 447, 46]. Statistical [660, 678, 156, 518, 599, 3, 609, 647, 394, 536, 70, 267, 186, 172, 285, 557, 546]. Statistics [605, 684, 434, 615, 669, 418, 479, 461, 515, 212, 333, 194, 328, 582, 91, 574, 67, 464, 224, 228, 337, 293, 31, 106, 347, 14, 652, 656, 181, 151, 592]. Status [648, 291, 190]. Stein [231, 357, 504]. Stein-rules [504]. step [416]. Stochastic [93, 146, 675, 81, 381, 267, 135, 159, 568, 27, 42, 129, 121, 416, 579, 585, 541, 466, 554]. stratified [47]. strict [527, 362]. strictly [329]. Strong [536, 667, 322, 62, 61, 456, 492, 580, 74, 248, 213, 100, 510]. Strongly [158, 437, 247, 234]. Structural [564]. Structure [434, 435, 574, 459, 404]. structured [560]. Structures [664, 201]. Student [539, 547]. Studentized [335, 364]. Study [224]. sub [533]. sub-ballistic [533]. subclass [303]. Subject [677]. subsampling [279]. suffering [367]. Sufficiency [189]. sum [125]. Summaries [679]. sums [340, 62, 576, 218, 253]. sup [320, 222, 576]. sup-functionals [576]. sup-norm [320, 222]. Superefficient [178]. superkernel [484]. supermartingale [587]. Supervised [514]. Supplement [253]. support [113, 206]. supported [358]. Supremum [642]. sure [325]. survey [570, 22, 13, 11]. survival [41]. switching [395]. Symmetric [664, 652, 656, 602, 497, 5, 55, 142]. symmetrization [69]. symmetrized [141]. symmetry [186, 174, 92, 465]. Systems [668, 225, 165, 554, 571].

T [434]. tables [465]. Tail [314, 672, 615, 665, 663, 547]. tailed [559, 279]. Tails [625]. tangents [180]. TAR [490]. tensor [425]. terms [554]. Terrell [671]. tessellations [29]. Test [653, 657, 623, 363, 610, 596, 136, 315, 487, 378, 387, 524, 332, 5, 333, 334, 328, 18, 442, 433, 92, 355, 383, 83, 298, 473]. Testing [43, 362, 406, 163, 630, 641, 230, 434, 278, 576, 432, 15, 644, 435, 319, 569, 527, 177, 200, 412, 520, 304, 472, 236, 225, 23, 17, 12, 24, 25, 126, 114, 306, 310, 368, 421, 478, 195, 566, 348, 155, 240, 404, 148]. Tests [618, 629, 31, 634, 376, 366, 66, 59, 36, 76, 455, 469, 536, 5, 103, 116, 550, 249, 193, 186, 255, 414, 256, 568, 9, 145, 329, 565, 274]. Theil [332]. their [200, 520, 125, 274]. theorem [35, 70, 280, 78, 510, 167, 197, 209]. Theorems [631, 468, 475, 212, 341, 228, 279, 61, 71, 53]. theoretic [270]. theoretical [401]. Theory [626, 605, 655, 272, 413, 381, 109, 48, 221, 246, 296, 506, 513, 10, 286, 292, 595, 181]. Third [32]. Third-order [32]. Three [328]. threshold [490]. tied [576]. tied-down [576]. Tikhonov [563]. Time [609, 645, 675, 558, 211, 200, 262, 169, 555, 215, 470, 518, 266, 458, 551, 483, 575, 120, 105, 121]. time-varying [558]. tomography [164, 354, 375]. Total [154, 584, 581]. tracking [558]. Transform [638, 478, 329]. transformation [412]. Transformed [684]. transient [3]. transition [214]. treatment [544]. treatments [245]. trend [178, 260]. Triangular [675, 228, 552]. trimmed [335, 364, 582]. Truncated [408, 681, 673, 615, 322, 579, 585, 495, 553, 531]. truncation [546]. Twice [637]. Two [596, 366, 199, 118, 645, 675, 77, 86, 111, 430, 103, 482, 472, 550, 544, 551, 337, 15, 561, 383]. two-component [430]. Two-sample [596, 337]. two-treatment [544]. two-way [103]. Type [618, 667, 515, 582, 100, 249, 306, 310, 37, 353, 63]. Types [647]. unbiased [55, 142, 252]. Uncertain [648]. uncorrelated [591]. unequal [442, 433]. unified [261, 590, 14]. Uniform [682, 667, 248, 415, 468, 475, 341, 68, 243, 51]. Uniform-in-Bandwidth [682]. uniformity [363, 83]. unimodal [197]. Units [652, 656]. univariate [503]. universal [100, 131]. universally [133]. unknown [339, 598, 457, 367, 112, 352, 251, 40]. unlinking [592]. unstable [108]. Upper [501, 506, 513, 210]. use [200]. using [596, 456, 545, 556].

V [655]. V. [149]. Validation [677]. value [272, 593, 559, 29, 224, 294]. Valued [663, 340, 566]. Values [674, 647, 662, 557]. Vapnik [342]. Varentropy [658]. variable [210, 328, 374]. Variables [401, 614, 664, 646, 226, 321, 125, 62, 591, 227, 223]. Variance [545, 562, 349, 382, 211, 115, 54, 412]. Variance-optimal [545]. Variate [655, 39]. variation [376, 581, 63, 154, 133]. Various [652, 656]. varying [558, 216, 594]. varying-coefficient [216]. vector [107]. Versions [173]. versus [243]. via [481, 608, 643, 289, 307, 519, 576, 432, 84, 675]. View [633, 184]. Vitality [669]. volatility [466]. Volterra [675].

walk [522, 533]. walks [584, 55, 142].

Warped [517]. Wasserstein [666]. Watson [249]. Watson-type [249]. wave [275]. Wavelet [112, 555, 58]. Wavelets [494]. Wavelets-based [494]. way [103, 465]. weak [602, 587, 30, 600, 356, 410]. weakly [75, 117, 234]. Weighted [424, 81, 527, 281, 89, 515, 599, 425, 576]. weighted-Gini [599]. weighting [523]. Weights [626]. white [339, 232]. Whitney [589]. whose [107]. width [166]. Wigner [354]. Wishart [655]. within [5]. without [37, 252, 546]. Wolfowitz [651]. work [176]. Woźniakowski [374].

Zero [660, 678, 303, 334]. Zero-Inflated [660, 678]. Zwet [635].

References

Belyaev:1992:KEH

[1] Yu. K. Belyaev. Kernel estimates of the hazard function based on incomplete data. Mathematical Methods of Statistics, 1:3–27, 1992. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kagan:1992:LPM

[2] Abram Kagan. Linearity of posterior mean for location parameter families when sample size equals one. Mathematical Methods of Statistics, 1:28–38, 1992. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kalashnikov:1992:SET

[3] V. V. Kalashnikov. Statistical estimates of transient periods for regenerative models by the coupling method. Mathematical Methods of Statistics, 1: 39–48, 1992. ISSN 1066-5307 (print), 1934-8045 (electronic).

Nevzorov:1992:CED

[4] V. B. Nevzorov. A characterization of exponential distributions by correlations between records. Mathematical Methods of Statistics, 1:49–54, 1992. ISSN 1066-5307 (print), 1934-8045 (electronic). See comments [149].

Chibisov:1992:AOC

[5] D. M. Chibisov. Asymptotic optimality of the chi-square test with large number of degrees of freedom within the class of symmetric tests. Mathematical Methods of Statistics, 1:55–82, 1992. ISSN 1066-5307 (print), 1934-8045 (electronic).

Spokoiny:1993:ODR

[6] V. G. Spokoiny. Optimal design of regression experiments with ARMAerrors. Mathematical Methods of Statistics, 2(1):1–17, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Goldenshluger:1993:ERP

[7] A. V. Goldenshluger and B. T. Polyak. Estimation of regression parameters with arbitrary noise. Mathematical Methods of Statistics, 2(1):18–29, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Piterbarg:1993:ADI

[8] V. I. Piterbarg and M. Ya. Penskaya. On asymptotic distribution of integrated squared error of an estimate of a component of a convolution. Mathematical Methods of Statistics, 2(1):30–41, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Luschgy:1993:APT

[9] H. Luschgy and A. L. Rukhin. Asymptotic properties of tests for a class of diffusion processes: optimality and adaptation. Mathematical Methods of Statistics, 2(1):42–51, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Levine:1993:ATP

[10] A. Levine and J. Liukkonen. Asymptotic theory for k-permutations of spacings. Mathematical Methods of Statistics, 2(1):52–71, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ivchenko:1993:PDSb

[11] G. I. Ivchenko and V. V. Lëvin. Processes in discrete schemes (a survey). I. Mathematical Methods of Statistics, 2 (1):72–84, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1993:AMHc

[12] Yu. I. Ingster. Asymptotically minimax hypothesis testing for nonparametric alternatives. I. Mathematical Methods of Statistics, 2(2):85–114, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic). See errata [25].

Ivchenko:1993:PDSa

[13] G. I. Ivchenko and V. V. Lëvin. Processes in discrete schemes (a survey). II. Mathematical Methods of Statistics, 2(2):115–129, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ruymgaart:1993:UAI

[14] Frits H. Ruymgaart. A unified approach to inversion problems in statistics. Mathematical Methods of Statistics, 2(2):130– 146, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Piterbarg:1993:THT

[15] V. I. Piterbarg and Yu. N. Tyurin. Testing for homogeneity of two multivariate samples: a Gaussian field on a sphere. Mathematical Methods of Statistics, 2(2):147–164, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bakirov:1993:EPB

[16] N. K. Bakirov. An extremal property of the binomial distribution. Mathematical Methods of Statistics, 2(2):165–170, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1993:AMHb

[17] Yu. I. Ingster. Asymptotically minimax hypothesis testing for nonparametric alternatives. II. Mathematical Methods of Statistics, 2(3):171–189, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic). See errata [24].

Larsson:1993:ADT

[18] Rolf Larsson. The asymptotic distribution of a test for the mean number of offspring in the branching process with immigration. Mathematical Methods of Statistics, 2(3):190–205, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Malinovskii:1993:AES

[19] V. K. Malinovskiĭ. Asymptotic expansions in sequential estimation of an autoregressive parameter. Mathematical Methods of Statistics, 2(3):206–227, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic). See erratum [34].

Rao:1993:ANL

[20] C. R. Rao and L. C. Zhao. Asymptotic normality of LAD estimator in censored regression models. Mathematical Methods of Statistics, 2(3):228–239, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Khmaladze:1993:PDL

[21] È. V. Khmaladze and Z. P. Tsigroshvili. On polynomial distributions with a large number of rare events. Mathematical Methods of Statistics, 2(3):240–247, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ivchenko:1993:EPD

[22] G. I. Ivchenko and V. V. Lëvin. Errata: “Processes in discrete schemes (a survey). II”. Mathematical Methods of Statistics, 2(3):248, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1993:AMHa

[23] Yu. I. Ingster. Asymptotically minimax hypothesis testing for nonparametric alternatives. III. Mathematical Methods of Statistics, 2(4):249–268, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1993:EAMa

[24] Yu. I. Ingster. Errata: “Asymptotically minimax hypothesis testing for nonparametric alternatives. II” [Math. Methods Statist. 2 (1993), no. 3, 171–189; MR1257983 (95g:62101)]. Mathematical Methods of Statistics, 2(4):268, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic). See [17].

Ingster:1993:EAMb

[25] Yu. I. Ingster. Errata: “Asymptotically minimax hypothesis testing for nonparametric alternatives. I” [Math. Methods Statist. 2 (1993), no. 2, 85– 114; MR1257978 (95g:62100)]. Mathematical Methods of Statistics, 2(4):268, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic). See [12].

Grigorev:1993:AEQ

[26] Yu. D. Grigor’ev and A. V. Ivanov. Asymptotic expansions for quadratic functionals of the least squares estimator of a nonlinear regression parameter. Mathematical Methods of Statistics, 2(4):269–294, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

LeBreton:1993:AAA

[27] Alain Le Breton. About the averaging approach in Gaussian schemes for stochastic approximation. Mathematical Methods of Statistics, 2(4):295–315, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Inglot:1993:ABS

[28] T. Inglot, W. C. M. Kallenberg, and T. Ledwina. Asymptotic behavior of some bilinear functionals of the empirical process. Mathematical Methods of Statistics, 2(4):316–336, 1993. ISSN 1066-5307 (print), 1934-8045 (electronic).

Heinrich:1994:NAS

[29] Lothar Heinrich. Normal approximation for some mean-value estimates of absolutely regular tessellations. Mathematical Methods of Statistics, 3(1):1–24, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gyorfi:1994:MHD

[30] L. Györfi, I. Vajda, and E. van der Meulen. Minimum Hellinger distance point estimates consistent under weak family regularity. Mathematical Methods of Statistics, 3(1):25–45, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rao:1994:TBS

[31] C. Radhakrishna Rao and Rahul Mukerjee. Tests based on score statistics: power properties and related results. Mathematical Methods of Statistics, 3 (1):46–61, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Zheng:1994:TOC

[32] Xiaodong Zheng. Third-order correct bootstrap calibrated confidence bounds for nonparametric mean. Mathematical Methods of Statistics, 3(1):62–75, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gill:1994:GCK

[33] Richard D. Gill. Glivenko–Cantelli for Kaplan–Meier. Mathematical Methods of Statistics, 3(1):76–87, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Malinovskii:1994:EAE

[34] V. K. Malinovskiĭ. Erratum: “Asymptotic expansions in sequential estimation of an autoregressive parameter”. Mathematical Methods of Statistics, 3(1):88, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic). See [19].

Coquet:1994:RCF

[35] F. Coquet, J. Mémin, and L. Vostrikova. Rate of convergence in the functional limit theorem for likelihood ratio processes. Mathematical Methods of Statistics, 3(2):89–113, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:1994:MET

[36] M. V. Boldin. On median estimates and tests in autoregressive models. Mathematical Methods of Statistics, 3(2):114– 129, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lahiri:1994:SOC

[37] Soumendra Nath Lahiri. On second order correctness of Efron’s bootstrap without Cramér-type conditions in linear regression models. Mathematical Methods of Statistics, 3(2):130–148, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Sun:1994:LBP

[38] Shan Sun and Martien C. A. van Zuijlen. Limiting behavior of the perturbed empirical distribution functions evaluated at a random point under dependence. Mathematical Methods of Statistics, 3(2):149–162, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gupta:1994:CMV

[39] A. K. Gupta and T. Varga. Characterization of matrix variate normality through conditional distributions. Mathematical Methods of Statistics, 3(2):163– 170, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Schick:1994:EER

[40] A. Schick. On efficient estimation in regression models with unknown scale functions. Mathematical Methods of Statistics, 3(3):171–212, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

vanderLaan:1994:MEE

[41] M. J. van der Laan. Modified EMestimator of the bivariate survival function. Mathematical Methods of Statistics, 3(3):213–243, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

LeBreton:1994:AEC

[42] A. Le Breton and A. A. Novikov. Averaging for estimating covariances in stochastic approximation. Mathematical Methods of Statistics, 3(3):244–266, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bakirov:1994:TNH

[43] N. K. Bakirov. Testing of nonparametric hypotheses for multivariate data. Mathematical Methods of Statistics, 3(3):267– 278, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:1994:NSP

[44] M. V. Boldin and Yu. N. Tyurin. On nonparametric sign procedures for autoregression models. Mathematical Methods of Statistics, 3(4):279–305, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Barbe:1994:LDM

[45] P. Barbe. Limiting distribution of the minimal spacing. Mathematical Methods of Statistics, 3(4):306–325, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Sherman:1994:ANG

[46] M. Sherman. Asymptotic normality for a general statistic from a random field. Mathematical Methods of Statistics, 3(4):326–345, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ivchenko:1994:MLE

[47] G. I. Ivchenko and S. A. Khonov. Maximum likelihood estimation for a stratified finite population. Mathematical Methods of Statistics, 3(4):346–361, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ferger:1994:ADT

[48] D. Ferger. Asymptotic distribution theory of change-point estimators and confidence intervals based on bootstrap approximation. Mathematical Methods of Statistics, 3(4):362–378, 1994. ISSN 1066-5307 (print), 1934-8045 (electronic).

Beirlant:1995:ANN

[49] J. Beirlant and D. M. Mason. On the asymptotic normality of L_{p}-norms of empirical functionals. Mathematical Methods of Statistics, 4(1):1–19, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

vandeGeer:1995:MSM

[50] S. van de Geer. The method of sieves and minimum contrast estimators. Mathematical Methods of Statistics, 4(1):20– 38, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Shi:1995:UBK

[51] Z. Shi. On the uniform Bahadur– Kiefer representation. Mathematical Methods of Statistics, 4(1):39–55, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Wu:1995:MKD

[52] C. O. Wu. Minimax kernel density estimators with length biased data. Mathematical Methods of Statistics, 4(1):56– 80, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rao:1995:CTE

[53] C. R. Rao and L. C. Zhao. Convergence theorems for empirical cumulative quantile regression functions. Mathematical Methods of Statistics, 4(1):81–91, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic). See correction [71].

Bshouty:1995:CVF

[54] D. Bshouty. On a characterization of variance functions of natural exponential families. Mathematical Methods of Statistics, 4(1):92–98, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mattner:1995:MUM

[55] L. Mattner. Mean unbiased medians, median unbiased means, and symmetric random walks. Mathematical Methods of Statistics, 4(1):99–105, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic). See addendum [142].

Cacoullos:1995:GCI

[56] T. Cacoullos and V. Papathanasiou. A generalization of covariance identity and related characterizations. Mathematical Methods of Statistics, 4(1):106–113, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:1995:LGA

[57] J. Pfanzagl. On local and global asymptotic normality. Mathematical Methods of Statistics, 4(2):115–136, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Neumann:1995:EWE

[58] M. H. Neumann and V. G. Spokoiny. On the efficiency of wavelet estimators under arbitrary error distributions. Mathematical Methods of Statistics, 4(2):137– 166, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bening:1995:FDOb

[59] V. E. Bening. A formula for deficiency: one sample L- and R-tests. I. Mathematical Methods of Statistics, 4(2):167–188, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Aerts:1995:BNP

[60] M. Aerts and N. Veraverbeke. Bootstrapping a nonparametric polytomous regression model. Mathematical Methods of Statistics, 4(2):189–200, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ralescu:1995:SAT

[61] S. S. Ralescu. Strong approximation theorems for integrated kernel quantiles. Mathematical Methods of Statistics, 4(2):201–215, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Liebscher:1995:SCS

[62] E. Liebscher. Strong convergence of sums of ϕ-mixing random variables. Mathematical Methods of Statistics, 4 (2):216–229, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Papadatos:1995:DVF

[63] N. Papadatos and V. Papathanasiou. Distance in variation and a Fishertype information. Mathematical Methods of Statistics, 4(2):230–237, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lepskii:1995:LAI

[64] O. V. Lepskiĭ and V. G. Spokoiny. Local adaptation to inhomogeneous smoothness: resolution level. Mathematical Methods of Statistics, 4(3):239–258, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Belitser:1995:MFE

[65] E. N. Belitser and B. Y. Levit. On minimax filtering over ellipsoids. Mathematical Methods of Statistics, 4(3):259–273, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bening:1995:FDOa

[66] V. E. Bening. A formula for deficiency: one-sample L- and R-tests. II. Mathematical Methods of Statistics, 4(3):274– 293, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ivchenko:1995:LDI

[67] G. I. Ivchenko and Sh. A. Mirakhmedov. Large deviations and intermediate efficiency of decomposable statistics in a multinomial scheme. Mathematical Methods of Statistics, 4(3):294–311, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Nagaev:1995:ABR

[68] A. V. Nagaev. Asymptotics of the Bayes risk in discrimination between normal and uniform location-scale families. Mathematical Methods of Statistics, 4(3):312–333, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mashayekhi:1995:SPM

[69] M. Mashayekhi. On symmetrization of product measures. Mathematical Methods of Statistics, 4(3):334–343, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dion:1995:CLT

[70] J.-P. Dion and N. M. Yanev. Central limit theorem for martingales in BGWR branching processes with some statistical applications. Mathematical Methods of Statistics, 4(3):344–358, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rao:1995:CCT

[71] C. R. Rao and L. C. Zhao. Correction: “Convergence theorems for empirical cumulative quantile regression functions”. Mathematical Methods of Statistics, 4 (3):359, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic). See [53].

Ghosal:1995:ABB

[72] Subhashis Ghosal and Tapas Samanta. Asymptotic behaviour of Bayes estimates and posterior distributions in multiparameter nonregular cases. Mathematical Methods of Statistics, 4(4):361– 388, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Tartakovsky:1995:APC

[73] A. Tartakovsky. Asymptotic properties of CUSUM and Shiryaev’s procedures for detecting a change in a nonhomogeneous Gaussian process. Mathematical Methods of Statistics, 4(4):389–404, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cheng:1995:OSC

[74] Ph. E. Cheng and Zhidong Bai. Optimal strong convergence rates in nonparametric regression. Mathematical Methods of Statistics, 4(4):405–420, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Konakov:1995:HLE

[75] V. D. Konakov and V. I. Piterbarg. High level excursions of Gaussian fields and the weakly optimal choice of the smoothing parameter. I. Mathematical Methods of Statistics, 4(4):421–434, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:1995:NST

[76] M. V. Boldin. On nonparametric sign tests in multiparameter autoregression. Mathematical Methods of Statistics, 4 (4):435–448, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Abadir:1995:JDT

[77] K. M. Abadir. The joint density of two functionals of Brownian motion. Mathematical Methods of Statistics, 4(4):449– 462, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic). See corrections [86, 111].

Janzura:1995:CLT

[78] M. Janžura and P. Lachout. A central limit theorem for stationary random fields. Mathematical Methods of Statistics, 4(4):463–471, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Karunamuni:1995:RGB

[79] R. J. Karunamuni and B. Schmuland. A robust generalized Bayes estimator of a multivariate normal mean. Mathematical Methods of Statistics, 4(4):472–482, 1995. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:1996:SOM

[80] G. K. Golubev and B. Y. Levit. On the second order minimax estimation of distribution functions. Mathematical Methods of Statistics, 5(1):1–31, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dippon:1996:WMP

[81] J. Dippon and J. Renz. Weighted means of processes in stochastic approximation. Mathematical Methods of Statistics, 5 (1):32–60, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Khasminskii:1996:PES

[82] R. Z. Khasminskii and B. V. Lazareva. Parameter estimation of a signal from linear indirect observations. Mathematical Methods of Statistics, 5(1):61–76, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Takacs:1996:TUC

[83] L. Takács. On a test for uniformity of a circular distribution. Mathematical Methods of Statistics, 5(1):77–98, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mukherjee:1996:REN

[84] K. Mukherjee. Robust estimation in nonlinear regression via minimum distance method. Mathematical Methods of Statistics, 5(1):99–112, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cohen:1996:LCR

[85] A. Cohen and H. B. Sackrowitz. Lower confidence regions for restricted parameter spaces. Mathematical Methods of Statistics, 5(1):113–123, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Abadir:1996:CJDa

[86] K. M. Abadir. Correction: “The joint density of two functionals of Brownian motion”. Mathematical Methods of Statistics, 5(1):124, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic). See [77].

Konev:1996:AMF

[87] V. V. Konev and S. M. Pergamenshchikov. On asymptotic minimaxity of fixed accuracy estimators for autoregression parameters. I. Stable process. Mathematical Methods of Statistics, 5(2):125– 153, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:1996:NSE

[88] M. V. Boldin. On nonparametric sign estimators in multiparameter autoregression. Mathematical Methods of Statistics, 5(2):154–172, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Borovkov:1996:ABW

[89] K. Borovkov, H. Dehling, and D. Pfeifer. On asymptotic behavior of weighted sample quantiles. Mathematical Methods of Statistics, 5(2):173–186, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Fountain:1996:SCE

[90] R. L. Fountain, J. P. Keating, and H. B. Maynard. The simultaneous comparison of estimators. Mathematical Methods of Statistics, 5(2):187–198, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Hallin:1996:ABC

[91] M. Hallin and K. Rifi. Asymptotic behavior of the characteristic function of simple serial rank statistics. Mathematical Methods of Statistics, 5(2):199–213, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Nikitin:1996:BHT

[92] Ya. Yu. Nikitin. On Baringhaus–Henze test for symmetry: Bahadur efficiency and local optimality for shift alternatives. Mathematical Methods of Statistics, 5(2):214–226, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Poix:1996:SAS

[93] J. Poix. Stochastic approximation for semimartingale error. Mathematical Methods of Statistics, 5(2):227–236, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kamiya:1996:BIB

[94] Hidehiko Kamiya. Borel isomorphism between the sample space and a product space. Mathematical Methods of Statistics, 5(2):237–243, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Raktoe:1996:CPS

[95] B. L. Raktoe and H. Pesotan. The critical problem of saturated s_{R}^{k−p} fractional factorial designs. Mathematical Methods of Statistics, 5(2):244–252, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Schipper:1996:ORC

[96] M. Schipper. Optimal rates and constants in L_{2}-minimax estimation of probability density functions. Mathematical Methods of Statistics, 5(3):253– 274, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Wu:1996:KSN

[97] Colin O. Wu. Kernel smoothing of the nonparametric maximum likelihood estimates for biased sampling models.

Mathematical Methods of Statistics, 5 (3):275–298, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Greenwood:1996:EES

[98] P. E. Greenwood and W. Wefelmeyer. Empirical estimators for semi-Markov processes. Mathematical Methods of Statistics, 5(3):299–315, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pensky:1996:EBE

[99] Marianna Pensky. Empirical Bayes estimation of a scale parameter. Mathematical Methods of Statistics, 5(3):316–331, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gyorfi:1996:SUC

[100] L. Györfi and H. Walk. On the strong universal consistency of a series type regression estimate. Mathematical Methods of Statistics, 5(3):332–342, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kukush:1996:ANE

[101] A. G. Kukush. Asymptotic normality of the estimator of an infinite-dimensional parameter in the model with a smooth regression function. Mathematical Methods of Statistics, 5(3):343–356, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:1996:AEE

[102] G. K. Golubev and B. Y. Levit. Asymptotically efficient estimation for analytic distributions. Mathematical Methods of Statistics, 5(3):357–368, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Clemmens:1996:LMP

[103] A. Clemmens and Z. Govindarajulu. Locally most powerful rank tests for random effects in two-way experiments. II. Mathematical Methods of Statistics, 5(3):369–382, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:1996:DFE

[104] G. K. Golubev and B. Y. Levit. Distribution function estimation: adaptive smoothing. Mathematical Methods of Statistics, 5(4):383–403, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Nze:1996:FET

[105] P. Ango Nze and P. Doukhan. Functional estimation for time series. I. Quadratic convergence properties. Mathematical Methods of Statistics, 5 (4):404–423, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rukhin:1996:CPE

[106] Andrew L. Rukhin. Change-point estimation: linear statistics and asymptotic Bayes risk. Mathematical Methods of Statistics, 5(4):424–442, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bischoff:1996:DWC

[107] Wolfgang Bischoff. On distribution whose conditional distributions are normal. A vector space approach. Mathematical Methods of Statistics, 5(4):443– 463, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kubilius:1996:AAL

[108] K. Kubilius and A. Račkauskas. On the asymptotic accuracy of least-squares estimators in nearly unstable AR(1) processes. Mathematical Methods of Statistics, 5(4):464–476, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dmitrienko:1996:RTR

[109] A. A. Dmitrienko and A. A. Vexler.

Renewal theory results for autoregressive processes. Mathematical Methods of Statistics, 5(4):477–490, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Keyes:1996:CSP

[110] T. K. Keyes and M. S. Levy. On consistency of some predicting densities for the multivariate linear model. Mathematical Methods of Statistics, 5(4):491–504, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic).

Abadir:1996:CJDb

[111] K. M. Abadir. Correction: “The joint density of two functionals of a Brownian motion” [Math. Methods Statist. 4 (1995), no. 4, 449–462; MR1372016 (97c:60200a)]. Mathematical Methods of Statistics, 5(4):505, 1996. ISSN 1066-5307 (print), 1934-8045 (electronic). See [77].

Juditsky:1997:WEA

[112] A. Juditsky. Wavelet estimators: adapting to unknown smoothness. Mathematical Methods of Statistics, 6(1):1–25, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gayraud:1997:EFD

[113] G. Gayraud. Estimation of functionals of density support. Mathematical Methods of Statistics, 6(1):26–46, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1997:SPH

[114] Yu. I. Ingster. Some problems of hypothesis testing leading to infinitely divisible distributions. Mathematical Methods of Statistics, 6(1):47–69, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic). See correction [126].

Boldin:1997:REE

[115] M. V. Boldin. On rank estimates and the empirical distribution function of residuals in autoregression with a possibly infinite variance. Mathematical Methods of Statistics, 6(1):70–91, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Govindarajulu:1997:CAD

[116] Z. Govindarajulu. A class of asymptotically distribution-free tests for equality of marginals in multivariate populations. Mathematical Methods of Statistics, 6(1):92–111, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Konakov:1997:HLE

[117] V. D. Konakov and V. I. Piterbarg. High level excursions of Gaussian fields and the weakly optimal choice of the smoothing parameter. II. Mathematical Methods of Statistics, 6(1):112–124, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Hopfner:1997:TCP

[118] R. Höpfner. Two comments on parameter estimation in stable processes. Mathematical Methods of Statistics, 6(1):125– 134, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Goldenshluger:1997:SAE

[119] A. Goldenshluger and A. Nemirovski.

On spatially adaptive estimation of nonparametric regression. Mathematical Methods of Statistics, 6(2):135–170, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Leblanc:1997:DEC

[120] Frédérique Leblanc. Density estimation for a class of continuous time processes. Mathematical Methods of Statistics, 6(2):171–199, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Piterbarg:1997:APP

[121] L. Piterbarg and B. Rozovskii. On asymptotic problems of parameter estimation in stochastic PDE’s: discrete time sampling. Mathematical Methods of Statistics, 6(2):200–223, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Li:1997:ORL

[122] Gang Li. Optimal rate local smoothing in a multiplicative intensity counting process model. Mathematical Methods of Statistics, 6(2):224–244, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lee:1997:LAM

[123] Stephen M. S. Lee. Local asymptotic minimax estimation of functionals of asymptotically normal sampling distributions. Mathematical Methods of Statistics, 6(2):245–257, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Angus:1997:MBK

[124] John E. Angus. A modified Bochner kernel lemma for density estimators. Mathematical Methods of Statistics, 6(2):258– 262, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kagan:1997:ISA

[125] A. Kagan, R. C. Laha, and V. Rohatgi. Independence of the sum and absolute difference of independent random variables does not imply their normality. Mathematical Methods of Statistics, 6(2):263–265, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1997:CSP

[126] Yu. I. Ingster. Correction: “Some problems of hypothesis testing leading to infinitely divisible distributions”. Mathematical Methods of Statistics, 6(2):266, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic). See [114].

Arcones:1997:GAB

[127] Miguel A. Arcones and David M. Mason. A general approach to Bahadur–Kiefer representations for M-estimators. Mathematical Methods of Statistics, 6(3):267– 292, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Scheffel:1997:RRL

[128] P. Scheffel and H. v. Weizsäcker. On risk rates and large deviations in finite Markov chain experiments. Mathematical Methods of Statistics, 6(3):293–312, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

LeBreton:1997:AFG

[129] Alain Le Breton. Averaging with feedback in Gaussian schemes for stochastic approximation. Mathematical Methods of Statistics, 6(3):313–331, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ghosal:1997:NAP

[130] Subhashis Ghosal. Normal approximation to the posterior distribution for generalized linear models with many covariates. Mathematical Methods of Statistics, 6(3):332–348, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kohler:1997:UCL

[131] Michael Kohler. On the universal consistency of a least squares spline regression estimator. Mathematical Methods of Statistics, 6(3):349–364, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Korostelev:1997:MLD

[132] Alexander Korostelev. Minimax large deviations risk in change-point problems. Mathematical Methods of Statistics, 6(3):365–374, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Viharos:1997:EER

[133] László Viharos. Estimators of the exponent of regular variation with universally normal asymptotic distributions. Mathematical Methods of Statistics, 6(3):375– 384, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Tracy:1997:EDS

[134] D. S. Tracy and S. S. Osahan. On the efficiency of determinant sampling. Mathematical Methods of Statistics, 6(3):385– 394, 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Huebner:1997:CAB

[135] M. Huebner. A characterization of asymptotic behaviour of maximum likelihood estimators for stochastic PDE’s. Mathematical Methods of Statistics, 6 (4):395–415 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Albers:1997:RTL

[136] W. Albers, W. C. M. Kallenberg, and G. D. Otten. Robust test limits. Mathematical Methods of Statistics, 6(4):416– 439 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Konakov:1997:SKD

[137] V. Konakov and E. Mammen. The shape of kernel density estimates in higher dimensions. Mathematical Methods of Statistics, 6(4):440–464 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:1997:ABA

[138] J. Pfanzagl. An asymptotic bound for the average length of confidence intervals. Mathematical Methods of Statistics, 6(4):465–475 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Leonov:1997:SOR

[139] Sergei L. Leonov. On the solution of an optimal recovery problem and its applications in nonparametric regression. Mathematical Methods of Statistics, 6 (4):476–490 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Tartakovsky:1997:MIR

[140] A. G. Tartakovsky. Minimax invariant regret solution to the N-sample slippage problem. Mathematical Methods of Statistics, 6(4):491–508 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Liu:1997:CRS

[141] Zhihui Liu. Convergence rates of symmetrized product measures. Mathematical Methods of Statistics, 6(4):509–516 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mattner:1997:APA

[142] L. Mattner. Acknowledgement of priority. Addendum to: “Mean unbiased medians, median unbiased means, and symmetric random walks” [Math. Methods Statist. 4 (1995), no. 1, 99–105; MR1324693 (96c:62029)]. Mathematical Methods of Statistics, 6(4):517 (1998), 1997. ISSN 1066-5307 (print), 1934-8045 (electronic). See [55].

Strasser:1998:PIE

[143] H. Strasser. Perturbation invariant estimates and incidental nuisance parameters. Mathematical Methods of Statistics, 7(1):1–26, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Konev:1998:AMF

[144] V. V. Konev and S. M. Pergamenshchikov. On asymptotic minimaxity of fixed accuracy estimators for autoregression parameters. II. Purely explosive process. Mathematical Methods of Statistics, 7(1):27–59, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Maglaperidze:1998:GFT

[145] N. O. Maglaperidze, Z. P. Tsigroshvili, and M. van Pul. Goodness-of-fit tests for parametric hypotheses on the distribution of point processes. Mathematical Methods of Statistics, 7(1):60–77, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Poix:1998:SAL

[146] J. Poix. Stochastic approximation for local martingale error and random gain process. Mathematical Methods of Statistics, 7(1):78–97, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Barbe:1998:LDP

[147] Ph. Barbe and M. Broniatowski. Large deviation principle for sequential rank processes. Mathematical Methods of Statistics, 7(1):98–110, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Volodin:1998:ANS

[148] I. N. Volodin and An. A. Novikov. Asymptotics of the necessary sample size in testing parametric hypotheses: dposterior approach. Mathematical Methods of Statistics, 7(1):111–121, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Szekely:1998:PVB

[149] G. J. Szekely and A. K. Gupta. On a paper of V. B. Nevzorov: “A characterization of exponential distributions by correlations between records” [Math. Methods Statist. 1 (1992), 49–54; MR1202799 (93m:62036)]. Mathematical Methods of Statistics, 7(1):122, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic). See [4].

Lepski:1998:AME

[150] O. V. Lepski and B. Y. Levit. Adaptive minimax estimation of infinitely differentiable functions. Mathematical Methods of Statistics, 7(2):123–156, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Suquet:1998:CPD

[151] Ch. Suquet and M.-C. Viano. Change point detection in dependent sequences: invariance principles for some quadratic statistics. Mathematical Methods of Statistics, 7(2):157–191, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:1998:AOF

[152] G. Golubev and R. Khasminskii.

Asymptotically optimal filtering for a hidden Markov model. Mathematical Methods of Statistics, 7(2):192–209, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Janaszewska:1998:JDS

[153] E. Janaszewska and A. V. Nagaev.

On the joint distribution of the shorth height and length. Mathematical Methods of Statistics, 7(2):210–229, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Papadatos:1998:TVD

[154] N. Papadatos and V. Papathanasiou.

Total variation distance and generalized covariance kernels. Mathematical Methods of Statistics, 7(2):230–244, 1998.

ISSN 1066-5307 (print), 1934-8045 (electronic).

Spokoiny:1998:ASA

[155] V. G. Spokoiny. Adaptive and spatially adaptive testing of a nonparametric hypothesis. Mathematical Methods of Statistics, 7(3):245–273, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bernstein:1998:SAR

[156] A. V. Bernstein. Statistical analysis of random polynomials. Mathematical Methods of Statistics, 7(3):274–295, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Chitashvili:1998:REP

[157] R. Chitashvili and N. Kordzakhia. Robust estimation of the parameter in a contaminated Markov model. Mathematical Methods of Statistics, 7(3):296– 318, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Fiorin:1998:SCF

[158] S. Fiorin. Strongly consistent functional estimation for a Cox regression model for counting processes. Mathematical Methods of Statistics, 7(3):319– 338, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Jankunas:1998:EPL

[159] A. Jankunas and R. Z. Khasminskii. Estimation of parameters of linear stochastic difference equations. Mathematical Methods of Statistics, 7(3):339–352, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pusz:1998:SCG

[160] J. Pusz. On some characterizations of the generalized inverse Gaussian distribution by regression with respect to residuals. Mathematical Methods of Statistics, 7(3):353–360, 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Koshkin:1998:NED

[161] G. M. Koshkin and V. A. Vasil’iev.

Nonparametric estimation of derivatives of a multivariate density from dependent observations. Mathematical Methods of Statistics, 7(4):361–400 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:1998:MDS

[162] Yu. I. Ingster. Minimax detection of a signal for l^{n}-balls. Mathematical Methods of Statistics, 7(4):401–428 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Conti:1998:TIL

[163] P. L. Conti and M. Scanu. Testing for independence in lattice distributions. Mathematical Methods of Statistics, 7 (4):429–444 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cavalier:1998:AEE

[164] L. Cavalier. Asymptotically efficient estimation in a problem related to tomography. Mathematical Methods of Statistics, 7(4):445–456 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kutoyants:1998:SED

[165] Yu. A. Kutoyants. Semiparametric estimation for dynamical systems with small noise. Mathematical Methods of Statistics, 7(4):457–465 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Basu:1998:FWC

[166] A. K. Basu and D. Bhattacharya. On fixed width confidence intervals associated with maximum likelihood estimation of parameter under dependent setup. Mathematical Methods of Statistics, 7(4):466–478 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Laha:1998:EDS

[167] R. G. Laha. An extension of the Darmois–Skitovich theorem. Mathematical Methods of Statistics, 7(4):479–483 (1999), 1998. ISSN 1066-5307 (print), 1934-8045 (electronic).

Menneteau:1999:FLI

[168] L. Menneteau. Functional law of the iterated logarithm for Kiefer processes. Mathematical Methods of Statistics, 8 (1):1–21, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bosq:1999:LTD

[169] D. Bosq and Yu. Davydov. Local time and density estimation in continuous time. Mathematical Methods of Statistics, 8(1):22–45, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Oudshoorn:1999:AEF

[170] C. G. M. Oudshoorn. Adaptive estimation of functions with jumps. Mathematical Methods of Statistics, 8(1):46– 68, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:1999:OLD

[171] J. Pfanzagl. On the optimality of limit distributions. Mathematical Methods of Statistics, 8(1):69–83, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). See erratum [196].

Kharin:1999:RMS

[172] Yu. Kharin. Robustness of multivariate statistical classification under nonparametric distortions of hypothetical distributions. Mathematical Methods of Statistics, 8(1):84–98, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Hille:1999:VCE

[173] J. Hille, D. Plachky, and J. Roters. Versions of conditional expectations depending continuously on parameters. Mathematical Methods of Statistics, 8(1):99–108, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mattner:1999:BSP

[174] L. Mattner. On Burdick’s symmetry problem. Mathematical Methods of Statistics, 8(1):109–118, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gupta:1999:ILM

[175] A. K. Gupta, G. J. Székely, and G. Zsigri. An inconsistent location MLE. Mathematical Methods of Statistics, 8 (1):119–120, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gotze:1999:WJP

[176] Friedrich Götze and Hartmut Milbrodt. The work of Johann Pfanzagl. Mathematical Methods of Statistics, 8(2):121– 141, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

Bening:1999:HOA

[177] V. E. Bening and D. M. Chibisov. Higher order asymptotic optimality in testing problems with nuisance parameters. I. Mathematical Methods of Statistics, 8(2):142–165, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

Beran:1999:SEM

[178] Rudolf Beran. Superefficient estimation of multivariate trend. Mathematical Methods of Statistics, 8(2):166–180, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

Greenwood:1999:BAE

[179] P. E. Greenwood and I. A. Ibragimov. Bahadur’s asymptotic efficiency and the LAN expansion. I. Lower bounds. Independent observations. Mathematical Methods of Statistics, 8(2):181–208, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

LeCam:1999:AT

[180] Lucien Le Cam. About tangents. Mathematical Methods of Statistics, 8(2):209– 219, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

Strasser:1999:ATP

[181] H. Strasser and Ch. Weber. The asymptotic theory of permutation statistics. Mathematical Methods of Statistics, 8 (2):220–250, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

vanderLaan:1999:ENN

[182] M. J. van der Laan and R. D. Gill.

Efficiency of NPMLE in nonparametric missing data models. Mathematical Methods of Statistics, 8(2):251–276, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

vanZwet:1999:RCE

[183] E. W. van Zwet and W. R. van Zwet. A remark on consistent estimation. Mathematical Methods of Statistics, 8(2):277– 284, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). In honor of the 70th birthday of Johann Pfanzagl.

Birge:1999:ICN

[184] Lucien Birgé. Interval censoring: a nonasymptotic point of view. Mathematical Methods of Statistics, 8(3):285–298, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gotze:1999:RBI

[185] F. Götze and B. A. Zalesky. Restoration of binary images for Bernoullian noise models. Mathematical Methods of Statistics, 8(3):299–319, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Janssen:1999:NST

[186] A. Janssen. Nonparametric symmetry tests for statistical functionals. Mathematical Methods of Statistics, 8(3):320– 343, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lepski:1999:ANE

[187] O. V. Lepski and B. Y. Levit. Adaptive nonparametric estimation of smooth multivariate functions. Mathematical Methods of Statistics, 8(3):344–370, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Liese:1999:SBP

[188] F. Liese and K. J. Miescke. Selection of the best population in models with nuisance parameters. Mathematical Methods of Statistics, 8(3):371–396, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mattner:1999:SEF

[189] L. Mattner. Sufficiency, exponential families, and algebraically independent numbers. Mathematical Methods of Statistics, 8(3):397–406, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Murphy:1999:CSR

[190] S. A. Murphy, A. W. van der Vaart, and J. A. Wellner. Current status regression. Mathematical Methods of Statistics, 8(3):407–425, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Schick:1999:EEI

[191] A. Schick and W. Wefelmeyer. Efficient estimation of invariant distributions of some semiparametric Markov chain models. Mathematical Methods of Statistics, 8(3):426–440, 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lepski:1999:HIA

[192] O. V. Lepski. How to improve the accuracy of estimation. Mathematical Methods of Statistics, 8(4):441–486 (2000), 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Inglot:1999:GIE

[193] T. Inglot. Generalized intermediate efficiency of goodness-of-fit tests. Mathematical Methods of Statistics, 8(4):487– 509 (2000), 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Fatalov:1999:LMC

[194] V. R. Fatalov. The Laplace method for computing exact asymptotics of distributions of integral statistics. Mathematical Methods of Statistics, 8(4):510–535 (2000), 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pouet:1999:TNH

[195] Christophe Pouet. On testing nonparametric hypotheses for analytic regression functions in Gaussian noise. Mathematical Methods of Statistics, 8(4):536–549 (2000), 1999. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:1999:EOL

[196] J. Pfanzagl. Erratum: “On the optimality of limit distributions” [Math. Methods Statist. 8 (1999), no. 1, 69– 83; MR1692715 (2000c:62079)]. Mathematical Methods of Statistics, 8(4):550 (2000), 1999. ISSN 1066-5307 (print), 1934-8045 (electronic). See [171].

Pfanzagl:2000:QCL

[197] J. Pfanzagl. Quasi-convex loss functions, unimodal distributions, and the convolution theorem. Mathematical Methods of Statistics, 9(1):1–18, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Tribouley:2000:AEI

[198] K. Tribouley. Adaptive estimation of integrated functionals. Mathematical Methods of Statistics, 9(1):19–38, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Butucea:2000:TAR

[199] C. Butucea. Two adaptive rates of convergence in pointwise density estimation. Mathematical Methods of Statistics, 9(1):39–64, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:2000:EPH

[200] M. V. Boldin. On empirical processes in heteroscedastic time series and their use for hypothesis testing and estimation. Mathematical Methods of Statistics, 9(1):65–89, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Plachky:2000:CPD

[201] D. Plachky and A. L. Rukhin. Characterization of probability distributions by independence structures. Mathematical Methods of Statistics, 9(1):90–105, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lababidi:2000:NES

[202] S. Lababidi. Nonparametric estimation of some functional with small noise diffusion processes. Mathematical Methods of Statistics, 9(1):106–116, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Efromovich:2000:SAE

[203] S. Efromovich. On sharp adaptive estimation of multivariate curves. Mathematical Methods of Statistics, 9(2):117– 139, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kim:2000:ESF

[204] Jae-Chun Kim and A. Korostelev. Estimation of smooth functionals in image models. Mathematical Methods of Statistics, 9(2):140–159, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:2000:SOM

[205] G. Golubev and W. Härdle. Second order minimax estimation in partial linear models. Mathematical Methods of Statistics, 9(2):160–175, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Potzelberger:2000:GQP

[206] K. Pötzelberger. The general quantization problem for distributions with regular support. Mathematical Methods of Statistics, 9(2):176–198, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Potzelberger:2000:CEQ

[207] K. Pötzelberger. Consistency of the empirical quantization error. Mathematical Methods of Statistics, 9(2):199–207, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Hadjicostas:2000:RCP

[208] P. Hadjicostas. Rate of convergence of the probability of non-existence of the MLE’s in simple logistic regression. Mathematical Methods of Statistics, 9(2):208–222, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Potzelberger:2000:CTI

[209] K. Pötzelberger, W. Schachermayer, and H. Strasser. The convolution theorem for infinite-dimensional parameter spaces. Mathematical Methods of Statistics, 9(3):223–236, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bischoff:2000:OLU

[210] W. Bischoff and M. Fichter. Optimal lower and upper bounds for the L_{p}-mean deviation of functions of a random variable. Mathematical Methods of Statistics, 9(3):237–269, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Blanke:2000:EAV

[211] D. Blanke and F. Merlevède. Estimation of the asymptotic variance of kernel density estimators for continuous time processes. Mathematical Methods of Statistics, 9(3):270–296, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Davydov:2000:FLT

[212] Yu. Davydov and V. Egorov. Functional limit theorems for induced order statistics. Mathematical Methods of Statistics, 9(3):297–313, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Einmahl:2000:SAS

[213] J. H. J. Einmahl and M. Geilen. A strong approximation of the shortt process. Mathematical Methods of Statistics, 9(3):314–322, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Clemencon:2000:AET

[214] S. J. M. Clémençon. Adaptive estimation of the transition density of a regular Markov chain. Mathematical Methods of Statistics, 9(4):323–357, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Camlong-Viot:2000:ATS

[215] Ch. Camlong-Viot, P. Sarda, and Ph.

Vieu. Additive time series: the kernel integration method. Mathematical Methods of Statistics, 9(4):358–375, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pons:2000:NEV

[216] Odile Pons. Nonparametric estimation in a varying-coefficient Cox model. Mathematical Methods of Statistics, 9 (4):376–398, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Einmahl:2000:SRE

[217] John H. J. Einmahl and Frits H. Ruymgaart. Some results for empirical processes of locally dependent arrays. Mathematical Methods of Statistics, 9(4):399– 414, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Novak:2000:SNS

[218] S. Y. Novak. On self-normalized sums.

Mathematical Methods of Statistics, 9 (4):415–436, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Yu:2000:APH

[219] Jun Yu and M. Ekström. Asymptotic properties of high order spacings under dependence assumptions. Mathematical Methods of Statistics, 9(4):437–448, 2000. ISSN 1066-5307 (print), 1934-8045 (electronic).

Holevo:2001:IQE

[220] A. S. Holevo. An invitation to quantum estimation. Mathematical Methods of Statistics, 10(1):1–5, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Holevo:2001:NAC

[221] A. S. Holevo. Noncommutative analogues of the Cramér-Rao inequality in the quantum measurement theory. Mathematical Methods of Statistics, 10 (1):6–22, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). See erratum [246].

Golubev:2001:AES

[222] Yu. Golubev, O. Lepski, and B. Levit. On adaptive estimation for the sup-norm losses. Mathematical Methods of Statistics, 10(1):23–37, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Worms:2001:MDSb

[223] J. Worms. Moderate deviations of some dependent variables. I. Martingales. Mathematical Methods of Statistics, 10(1):38–72, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lachal:2001:SSN

[224] Aimé Lachal. Study of some new integrated statistics: computation of Bahadur efficiency, relation with nonstandard boundary value problems. Mathematical Methods of Statistics, 10 (1):73–104, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Iacus:2001:SHT

[225] S. M. Iacus and Yu. A. Kutoyants. Semiparametric hypotheses testing for dynamical systems with small noise. Mathematical Methods of Statistics, 10(1): 105–120, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Comte:2001:SEA

[226] F. Comte and M.-L. Taupin. Semiparametric estimation in the (auto)- regressive β-mixing model with errorsin-variables. Mathematical Methods of Statistics, 10(2):121–160, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Worms:2001:MDSa

[227] J. Worms. Moderate deviations of some dependent variables. II. Some kernel estimators. Mathematical Methods of Statistics, 10(2):161–193, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Liebscher:2001:CLT

[228] Eckhard Liebscher. Central limit theorems for α-mixing triangular arrays with applications to nonparametric statistics. Mathematical Methods of Statistics, 10 (2):194–214, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lemdani:2001:RDK

[229] M. Lemdani and E. Ould-Said. Relative deficiency of the Kaplan–Meier estimator with respect to a smoothed estimator. Mathematical Methods of Statistics, 10(2):215–231, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Henze:2001:TEA

[230] N. Henze and B. Klar. Testing exponentiality against the L-class of life distributions. Mathematical Methods of Statistics, 10(2):232–246, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cavalier:2001:PBS

[231] L. Cavalier and A. B. Tsybakov. Penalized blockwise Stein’s method, monotone oracles and sharp adaptive estimation. Mathematical Methods of Statistics, 10(3):247–282, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Delattre:2001:PBM

[232] S. Delattre and M. Hoffmann. The Pinsker bound in mixed Gaussian white noise. Mathematical Methods of Statistics, 10(3):283–315, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Galtchouk:2001:SNA

[233] L. Galtchouk and S. Pergamenshchikov. Sequential nonparametric adaptive estimation of the drift coefficient in diffusion processes. Mathematical Methods of Statistics, 10(3):316–330, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Soulier:2001:AES

[234] Ph. Soulier. Adaptive estimation of the spectral density of a weakly or strongly dependent Gaussian process. Mathematical Methods of Statistics, 10(3):331– 354, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

vandeGeer:2001:LSE

[235] Sara van de Geer. Least squares estimation with complexity penalties. Mathematical Methods of Statistics, 10(3):355– 374, 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Gayraud:2001:MTC

[236] G. Gayraud and Ch. Pouet. Minimax testing composite null hypotheses in the discrete regression scheme. Mathematical Methods of Statistics, 10(4): 375–394 (2002), 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Ingster:2001:ADS

[237] Yu. I. Ingster. Adaptive detection of a signal of growing dimension. I. Mathematical Methods of Statistics, 10(4): 395–421 (2002), 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Juditsky:2001:EAN

[238] A. Juditsky and O. Lepski. Evaluation of the accuracy of nonparametric estimators. Mathematical Methods of Statistics, 10(4):422–445 (2002), 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Malyutov:2001:SOO

[239] M. B. Malyutov and I. I. Tsitovich. Second order optimal sequential discrimination between Markov chains. Mathematical Methods of Statistics, 10(4): 446–464 (2002), 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Spokoiny:2001:DDT

[240] V. Spokoiny. Data-driven testing the fit of linear models. Mathematical Methods of Statistics, 10(4):465–497 (2002), 2001. ISSN 1066-5307 (print), 1934-8045 (electronic). Meeting on Mathematical Statistics (Marseille, 10–15 December 2000).

Grama:2002:AEN

[241] I. Grama and M. Nussbaum. Asymptotic equivalence for nonparametric regression. Mathematical Methods of Statistics, 11(1):1–36, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:2002:ADS

[242] Yu. I. Ingster. Adaptive detection of a signal of growing dimension. II. Mathematical Methods of Statistics, 11(1):37– 68, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:2002:AOE

[243] J. Pfanzagl. Asymptotic optimality of estimator sequences: “pointwise” versus “locally uniform”. Mathematical Methods of Statistics, 11(1):69–97, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:2002:ASP

[244] G. Golubev and W. Härdle. On adaptive smoothing in partial linear models. Mathematical Methods of Statistics, 11(1):98–117, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bathke:2002:ALN

[245] Arne Bathke. ANOVA for a large number of treatments. Mathematical Methods of Statistics, 11(1):118–132, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Holevo:2002:ENA

[246] A. S. Holevo. Erratum: “Noncommutative analogues of the Cramér–Rao inequality in the quantum measurement theory” [Math. Methods Statist. 10 (2001), no. 1, 6–22; MR1841806 (2002e:81014)]. Mathematical Methods of Statistics, 11(1):133, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic). See [221].

Dehay:2002:SES

[247] D. Dehay and H. L. Hurd. Spectral estimation for strongly periodically correlated random fields defined on R^{2}. Mathematical Methods of Statistics, 11 (2):135–151, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Du:2002:USR

[248] Yunling Du and M. G. Akritas. Uniform strong representation of the conditional Kaplan–Meier process. Mathematical Methods of Statistics, 11(2):152– 182, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Henze:2002:WTG

[249] N. Henze and Ya. Yu. Nikitin. Watsontype goodness-of-fit tests based on the integrated empirical process. Mathematical Methods of Statistics, 11(2):183– 202, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Morel:2002:HIP

[250] B. Morel and Ch. Suquet. Hilbertian invariance principles for the empirical process under association. Mathematical Methods of Statistics, 11(2):203–220, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Muller:2002:EMC

[251] U. U. Müller and W. Wefelmeyer. Estimators for models with constraints involving unknown parameters. Mathematical Methods of Statistics, 11(2):221– 235, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Potzelberger:2002:AUQ

[252] K. Pötzelberger. Admissible unbiased quantizations: distributions without linear components. Mathematical Methods of Statistics, 11(2):236–255, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Novak:2002:SNS

[253] S. Y. Novak. Self-normalized sums. Supplement. Mathematical Methods of Statistics, 11(2):256–258, 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dette:2002:ODF

[254] H. Dette and V. B. Melas. E-optimal designs in Fourier regression models on a partial circle. Mathematical Methods of Statistics, 11(3):259–296 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Janssen:2002:HBA

[255] A. Janssen and J. Rahnenführer. A hazard-based approach to dependence tests for bivariate censored models. Mathematical Methods of Statistics, 11 (3):297–322 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kallenberg:2002:PDD

[256] W. C. M. Kallenberg. The penalty in data driven Neyman’s tests. Mathematical Methods of Statistics, 11(3):323–340 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rackauskas:2002:HCM

[257] A. Račkauskas and Ch. Suquet. Hölder convergence of multivariate empirical characteristic functions. Mathematical Methods of Statistics, 11(3):341–357 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Schick:2002:EEI

[258] A. Schick and W. Wefelmeyer. Efficient estimation in invertible linear processes. Mathematical Methods of Statistics, 11 (3):358–379 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bosq:2002:EAO

[259] D. Bosq. Estimation of the autocorrelation operator and prediction for infinite-dimensional autoregressive processes. Mathematical Methods of Statistics, 11(4):381–401 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dalalyan:2002:AET

[260] A. S. Dalalyan and Yu. A. Kutoyants. Asymptotically efficient trend coefficient estimation for ergodic diffusion. Mathematical Methods of Statistics, 11(4): 402–427 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dette:2002:UAE

[261] H. Dette and Yu. Grigoriev. A unified asymptotic expansion for distributions of quadratic functionals in nonlinear regression models. Mathematical Methods of Statistics, 11(4):428–452 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:2002:SRE

[262] M. V. Boldin. On sequential residual empirical processes in heteroscedastic time series. Mathematical Methods of Statistics, 11(4):453–464 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dumbgen:2002:ARC

[263] L. Dümbgen and V. R. Fatalov. Asymptotics of the rate of convergence for nonparametric density estimators: a new approach based on the Laplace method. Mathematical Methods of Statistics, 11 (4):465–476 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:2002:DLR

[264] J. Pfanzagl. On distinguished LANrepresentations. Mathematical Methods of Statistics, 11(4):477–488 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic). See erratum [324].

Balakrishnan:2002:CPA

[265] N. Balakrishnan and S. V. Malov. Characterizations of proportional Archimedean copulas. Mathematical Methods of Statistics, 11(4):489–495 (2003), 2002. ISSN 1066-5307 (print), 1934-8045 (electronic).

Guillou:2003:SOP

[266] A. Guillou and F. Merlevède. Secondorder properties of the blocks of blocks bootstrap for density estimators for continuous time processes. Mathematical Methods of Statistics, 12(1):1–30, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gushchin:2003:PSM

[267] A. A. Gushchin and Uwe Küchler. On parametric statistical models for stationary solutions of affine stochastic delay differential equations. Mathematical Methods of Statistics, 12(1):31–61, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Martinez:2003:AEA

[268] L. M. Artiles Martínez and B. Y. Levit.

Adaptive estimation of analytic functions on an interval. Mathematical Methods of Statistics, 12(1):62–94, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:2003:SRA

[269] G. Golubev and B. Levit. Sequential recovery of analytic periodic edges in binary image models. Mathematical Methods of Statistics, 12(1):95–115, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Chen:2003:ITA

[270] Jiahua Chen and A. K. Gupta. Information-theoretic approach for detecting change in the parameters of a normal model. Mathematical Methods of Statistics, 12(1):116–130, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Belitser:2003:EBA

[271] E. Belitser and B. Levit. On the empirical Bayes approach to adaptive filtering. Mathematical Methods of Statistics, 12(2):131–154, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Alves:2003:EPC

[272] M. I. Fraga Alves, L. de Haan, and Tao Lin. Estimation of the parameter controlling the speed of convergence in extreme value theory. Mathematical Methods of Statistics, 12(2):155–176, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Louani:2003:LDR

[273] D. Louani. Large deviations results for orthogonal series density estimators and some applications. Mathematical Methods of Statistics, 12(2):177–196, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Stepanova:2003:MRT

[274] N. A. Stepanova. Multivariate rank tests for independence and their asymptotic efficiency. Mathematical Methods of Statistics, 12(2):197–217, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Chow:2003:OIA

[275] P. L. Chow, R. Z. Khasminskii, and A. I.

Ovseevich. Optimal input and asymptotically efficient estimation in some wave equations. Mathematical Methods of Statistics, 12(2):218–230, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Vyazilov:2003:EPR

[276] A. E. Vyazilov. Empirical processes and robust estimation of parameters of the GARCH model. Mathematical Methods of Statistics, 12(2):231–245, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:2003:MNS

[277] Yu. Ingster and O. Lepski. Multichannel nonparametric signal detection. Mathematical Methods of Statistics, 12(3):247– 275, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Maller:2003:TIH

[278] R. A. Maller and Xian Zhou. Testing for individual heterogeneity in parametric models for event-history data. Mathematical Methods of Statistics, 12(3):276– 304, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

McElroy:2003:LTH

[279] T. McElroy and D. N. Politis. Limit theorems for heavy-tailed random fields with subsampling applications. Mathematical Methods of Statistics, 12(3):305– 328, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Haye:2003:NLT

[280] M. Ould Haye and A. Philippe. A noncentral limit theorem for the empirical process of linear sequences with seasonal long memory. Mathematical Methods of Statistics, 12(3):329–357, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Biedermann:2003:NMB

[281] S. Biedermann and H. Dette. A note on maximin and Bayesian D-optimal designs in weighted polynomial regression. Mathematical Methods of Statistics, 12(3):358–370, 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Biau:2003:SKD

[282] G. Biau. Spatial kernel density estimation. Mathematical Methods of Statistics, 12(4):371–390 (2004), 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Broniatowski:2003:EKL

[283] M. Broniatowski. Estimation of the Kullback–Leibler divergence. Mathematical Methods of Statistics, 12(4):391– 409 (2004), 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Juditsky:2003:NCS

[284] A. Juditsky and S. Lambert-Lacroix. Nonparametric confidence set estimation. Mathematical Methods of Statistics, 12(4):410–428 (2004), 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Klemela:2003:ORS

[285] Jussi Klemelä. Optimal recovery and statistical estimation in L_{p} Sobolev classes. Mathematical Methods of Statistics, 12(4):429–453 (2004), 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Liese:2003:GAT

[286] F. Liese and I. Vajda. A general asymptotic theory of M-estimators. I. Mathematical Methods of Statistics, 12(4): 454–477 (2004), 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Mnatsakanov:2003:SPM

[287] R. Mnatsakanov and F. H. Ruymgaart.

Some properties of moment-empirical CDF’s with application to some inverse estimation problems. Mathematical Methods of Statistics, 12(4):478–495 (2004), 2003. ISSN 1066-5307 (print), 1934-8045 (electronic).

Borovkov:2004:AOE

[288] A. A. Borovkov and Yu. Yu. Linke.

Asymptotically optimal estimates in the smooth change-point problem. Mathematical Methods of Statistics, 13(1):1– 24, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Galtchouk:2004:NSE

[289] L. Galtchouk and S. Pergamenshchikov.

Nonparametric sequential estimation of the drift in diffusion processes via model selection. Mathematical Methods of Statistics, 13(1):25–49, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Norberg:2004:FEI

[290] E. Norberg. On filtered experiments and the information contained in likelihood ratios. Mathematical Methods of Statistics, 13(1):50–68, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dumbgen:2004:CCR

[291] L. Dümbgen, S. Freitag, and G. Jongbloed. Consistency of concave regression with an application to current-status data. Mathematical Methods of Statistics, 13(1):69–81, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Liese:2004:GAT

[292] F. Liese and I. Vajda. A general asymptotic theory of M-estimators. II. Mathematical Methods of Statistics, 13(1):82– 95, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pycke:2004:BLA

[293] J.-R. Pycke. Bahadur local asymptotic optimality for generalizations of the Cramér-von Mises and Anderson– Darling statistics. Mathematical Methods of Statistics, 13(1):96–107, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Volodin:2004:PCV

[294] I. N. Volodin and S. V. Simushkin. Dposterior concept of p-value. Mathematical Methods of Statistics, 13(1):108–121, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Blanke:2004:SPA

[295] D. Blanke. Sample paths adaptive density estimation. Mathematical Methods of Statistics, 13(2):123–152, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ivanov:2004:ATN

[296] A. V. Ivanov and N. N. Leonenko. Asymptotic theory of nonlinear regression with long-range dependence. Mathematical Methods of Statistics, 13(2): 153–178, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Joutard:2004:LDE

[297] C. Joutard. Large deviations for M-estimators. Mathematical Methods of Statistics, 13(2):179–200, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Yode:2004:AMT

[298] A. F. Yode. Asymptotically minimax test of independence. Mathematical Methods of Statistics, 13(2):201–234, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kharin:2004:DAS

[299] Yu. Kharin and A. Kostevich. Discriminant analysis of stationary finite Markov chains. Mathematical Methods of Statistics, 13(2):235–252, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Levit:2004:EEM

[300] B. Levit and N. Stepanova. Efficient estimation of multivariate analytic functions in cube-like domains. Mathematical Methods of Statistics, 13(3):253–281, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Matias:2004:MEL

[301] C. Matias and M.-L. Taupin. Minimax estimation of linear functionals in the convolution model. Mathematical Methods of Statistics, 13(3):282–328, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Sorokin:2004:MDE

[302] A. A. Sorokin. Minimum distance estimates in the ARCH model. Mathematical Methods of Statistics, 13(3):329–355, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bar-Lev:2004:ZRC

[303] S. K. Bar-Lev, D. Bshouty, and F. A. Van der Duyn Schouten. Zero regression characterizations of natural exponential families generated by Lévy stable laws — a complementary subclass. Mathematical Methods of Statistics, 13(3):356– 367, 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gaffke:2004:NOS

[304] N. Gaffke. Nonparametric one-sided testing for the mean and related extremum problems. Mathematical Methods of Statistics, 13(4):369–391 (2005), 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:2004:OAA

[305] Y. Golubev and B. Levit. An oracle approach to adaptive estimation of linear functionals in a Gaussian model. Mathematical Methods of Statistics, 13 (4):392–408 (2005), 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:2004:NHT

[306] Yu. I. Ingster and I. A. Suslina. Nonparametric hypothesis testing for small type I errors. I. Mathematical Methods of Statistics, 13(4):409–459 (2005), 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kohler:2004:ESR

[307] M. Kohler. Estimation of smooth regression functions from stationary and ergodic observations via least squares. Mathematical Methods of Statistics, 13 (4):460–471 (2005), 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Schreiber:2004:LDP

[308] T. Schreiber. Large deviation principle for generalized quantile functions. Mathematical Methods of Statistics, 13 (4):472–483 (2005), 2004. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gaiffas:2005:CRP

[309] S. Gaïffas. Convergence rates for pointwise curve estimation with a degenerate design. Mathematical Methods of Statistics, 14(1):1–27, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:2005:NHT

[310] Yu. I. Ingster and I. A. Suslina. Nonparametric hypothesis testing for small type I errors. II. Mathematical Methods of Statistics, 14(1):28–52, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kukush:2005:CDEb

[311] A. Kukush and H. Schneeweiss. Comparing different estimators in a nonlinear measurement error model. I. Mathematical Methods of Statistics, 14(1):53–79, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Neumeyer:2005:NOS

[312] N. Neumeyer and H. Dette. A note on one-sided nonparametric analysis of covariance by ranking residuals. Mathematical Methods of Statistics, 14(1):80– 104, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ruckdeschel:2005:OOS

[313] P. Ruckdeschel. Optimally one-sided bounded influence curves. Mathematical Methods of Statistics, 14(1):105–131, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Albers:2005:TBE

[314] W. Albers and W. C. M. Kallenberg. Tail behavior of the empirical distribution function of convolutions. Mathematical Methods of Statistics, 14(2):133– 162, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Arcones:2005:BEL

[315] M. A. Arcones. Bahadur efficiency of the likelihood ratio test. Mathematical Methods of Statistics, 14(2):163–179, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ehm:2005:RKE

[316] Werner Ehm. On the risk of kernel estimators in a density reconstruction problem. Mathematical Methods of Statistics, 14(2):180–202, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Kukush:2005:CDEa

[317] A. Kukush and H. Schneeweiss. Comparing different estimators in a nonlinear measurement error model. II. Mathematical Methods of Statistics, 14(2):203– 223, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Matsui:2005:FIM

[318] M. Matsui. Fisher information matrix of general stable distributions close to the normal distribution. Mathematical Methods of Statistics, 14(2):224–251, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Abramovich:2005:LFH

[319] F. Abramovich and R. Heller. Local functional hypothesis testing. Mathematical Methods of Statistics, 14(3):253– 266, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Bertin:2005:SAE

[320] K. Bertin. Sharp adaptive estimation in sup-norm for d-dimensional Hölder classes. Mathematical Methods of Statistics, 14(3):267–298, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Ingster:2005:EDS

[321] Yu. Ingster and I. Suslina. On estimation and detection of smooth functions of many variables. Mathematical Methods of Statistics, 14(3):299–331, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Lemdani:2005:SRQ

[322] M. Lemdani, E. Ould-Saïd, and N. Poulin. Strong representation of the quantile function for left truncated and dependent data. Mathematical Methods of Statistics, 14(3):332–345, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rivoirard:2005:BMS

[323] V. Rivoirard. Bayesian modeling of sparse sequences and maxisets for Bayes rules. Mathematical Methods of Statistics, 14(3):346–376, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Pfanzagl:2005:EDL

[324] J. Pfanzagl. Erratum: “On distinguished LAN-representations” [Math. Methods Statist. 11 (2002), no. 4, 477– 488 (2003); MR1979746]. Mathematical Methods of Statistics, 14(3):377–378, 2005. ISSN 1066-5307 (print), 1934-8045 (electronic). See [264].

Berthet:2005:ASA

[325] P. Berthet and C. El-Nouty. Almost sure asymptotic behaviour of the shorth estimators. Mathematical Methods of Statistics, 14(4):379–403 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Borovkov:2005:CPP

[326] A. A. Borovkov and Yu. Yu. Linke. Change-point problem for large samples and incomplete information on distributions. Mathematical Methods of Statistics, 14(4):404–430 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cao:2005:OIR

[327] Y. Cao and Y. Golubev. On oracle inequalities related to a polynomial fitting. Mathematical Methods of Statistics, 14 (4):431–450 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gaffke:2005:TTS

[328] N. Gaffke. Three test statistics for a nonparametric one-sided hypothesis on the mean of a nonnegative variable. Mathematical Methods of Statistics, 14(4):

451–467 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Meintanis:2005:OTS

[329] S. Meintanis. Omnibus tests for strictly positive stable laws based on the empirical Laplace transform. Mathematical Methods of Statistics, 14(4):468–478 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Meister:2005:NES

[330] A. Meister. Non-estimability in spite of identifiability in density deconvolution. Mathematical Methods of Statistics, 14 (4):479–487 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Yu:2005:NCC

[331] Zhuo Yu and M. van der Laan. A note on the construction of counterfactuals and G-computation formula. Mathematical Methods of Statistics, 14(4):488–499 (2006), 2005. ISSN 1066-5307 (print), 1934-8045 (electronic).

Busarova:2006:MTE

[332] D. Busarova, Y. Tyurin, J. Möttönen, and H. Oja. Multivariate Theil estimator with the corresponding test. Mathematical Methods of Statistics, 15(1):1– 19, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dumbgen:2006:LDM

[333] L. Dümbgen, V. I. Piterbarg, and D. Zholud. On the limit distribution of multiscale test statistics for nonparametric curve estimation. Mathematical Methods of Statistics, 15(1):20–25, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Durot:2006:ATZ

[334] C. Durot and Y. Rozenholc. An adaptive test for zero mean. Mathematical Methods of Statistics, 15(1):26–60, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gribkova:2006:EEE

[335] N. V. Gribkova and R. Helmers. The empirical Edgeworth expansion for a Studentized trimmed mean. Mathematical Methods of Statistics, 15(1):61–87, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Korostelev:2006:EJP

[336] A. Korostelev and G. Yin. Estimation of jump points in high-dimensional diffusion modulated by a hidden Markov chain. Mathematical Methods of Statistics, 15(1):88–102, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Nikitin:2006:LDN

[337] Ya. Nikitin and E. Ponikarov. On large deviations of non-degenerate two-sample U- and V -statistics with applications to Bahadur efficiency. Mathematical Methods of Statistics, 15(1):103–122, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Autin:2006:MDE

[338] F. Autin. Maxiset for density estimation on R. Mathematical Methods of Statistics, 15(2):123–145, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Castillo:2006:EAE

[339] I. Castillo, C. Lévy-Leduc, and C. Matias. Exact adaptive estimation of the shape of a periodic function with unknown period corrupted by white noise. Mathematical Methods of Statistics, 15 (2):146–175, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Dedecker:2006:IPS

[340] J. Dedecker and F. Merlevède. Inequalities for partial sums of Hilbertvalued dependent sequences and applications. Mathematical Methods of Statistics, 15(2):176–206, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Hamadouche:2006:OHF

[341] D. Hamadouche and Ch. Suquet. Optimal Hölderian functional central limit theorems for uniform empirical and quantile processes. Mathematical Methods of Statistics, 15(2):207–223, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Khmaladze:2006:CSPb

[342] E. Khmaladze, R. Mnatsakanov, and N. Toronjadze. The change-set problem for Vapnik–Červonenkis classes. Mathematical Methods of Statistics, 15(2):224– 231, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Brunel:2006:ANR

[343] E. Brunel and F. Comte. Adaptive nonparametric regression estimation in presence of right censoring. Mathematical Methods of Statistics, 15(3):233–255, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cadre:2006:PBM

[344] B. Cadre and I. Larramendy-Valverde.

Prediction based on an L_{1}-method in the nonlinear autoregressive model. Mathematical Methods of Statistics, 15(3):256– 268, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Gupta:2006:IPP

[345] A. K. Gupta, T. Nguyen, and L. Pardo. Inference procedures for polytomous logistic regression models based on ϕdivergence measures. Mathematical Methods of Statistics, 15(3):269–288, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Khmaladze:2006:CSPa

[346] E. Khmaladze, R. Mnatsakanov, and N. Toronjadze. The change set problem and local covering numbers. Mathematical Methods of Statistics, 15(3):289–308, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Rukhin:2006:ACS

[347] A. L. Rukhin. Association characteristics in spatial statistics. Mathematical Methods of Statistics, 15(3):309–326, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Sorokin:2006:PET

[348] A. A. Sorokin. On parameter estimation and testing for dimension in ARCH(p) model. Mathematical Methods of Statistics, 15(3):327–348, 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Autin:2006:LVG

[349] F. Autin, D. Picard, and V. Rivoirard. Large variance Gaussian priors in Bayesian nonparametric estimation: a maxiset approach. Mathematical Methods of Statistics, 15(4):349–373 (2007), 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Neumann:2006:AME

[350] M. H. Neumann and T. L. Thorarinsdottir. Asymptotic minimax estimation in nonparametric autoregression. Mathematical Methods of Statistics, 15(4): 374–397 (2007), 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Cao:2006:OIR

[351] Y. Cao and Y. Golubev. On oracle inequalities related to smoothing splines. Mathematical Methods of Statistics, 15 (4):398–414 (2007), 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Marteau:2006:RIP

[352] C. Marteau. Regularization of inverse problems with unknown operator. Mathematical Methods of Statistics, 15(4): 415–443 (2007), 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Meintanis:2006:EMT

[353] S. G. Meintanis. Efficient moment-type estimation in exponentiated laws. Mathematical Methods of Statistics, 15(4): 444–455 (2007), 2006. ISSN 1066-5307 (print), 1934-8045 (electronic).

Guta:2007:MEW

[354] M. Guta and L. Artiles. Minimax estimation of the Wigner function in quantum homodyne tomography with ideal detectors. Mathematical Methods of Statistics, 16(1):1–15, March 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707010012.

Nikitin:2007:LTE

[355] Ya. Yu. Nikitin and A. V. Tchirina. Lilliefors test for exponentiality: Large deviations, asymptotic efficiency, and conditions of local optimality. Mathematical Methods of Statistics, 16(1):16–24, March 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707010024.

Prieur:2007:CPE

[356] C. Prieur. Change point estimation by local linear smoothing under a weak dependence condition. Mathematical Methods of Statistics, 16(1):25–41, March 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530707010036.

Dalalyan:2007:SSS

[357] A. S. Dalalyan. Stein shrinkage and second-order efficiency for semiparametric estimation of the shift. Mathematical Methods of Statistics, 16(1):42–62, March 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530707010048.

Meister:2007:DCS

[358] A. Meister. Deconvolving compactly supported densities. Mathematical Methods of Statistics, 16 (1):63–76, March 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S106653070701005X.

Anonymous:2007:HCa

[359] Anonymous. Help & contacts. Mathematical Methods of Statistics, 16(1):??, March 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Balabdaoui:2007:CEC

[360] F. Balabdaoui. Consistent estimation of a convex density at the origin. Mathematical Methods of Statistics, 16(2):77–95, June 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707020019.

Bar-Lev:2007:MCC

[361] S. K. Bar-Lev. Methods of constructing characterizations by constancy of regression on the sample mean and related problems for NEF’s. Mathematical Methods of Statistics, 16(2):96–109, June 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707020020.

Birke:2007:TSM

[362] M. Birke and H. Dette. Testing strict monotonicity in nonparametric regression. Mathematical Methods of Statistics, 16(2):110–123, June 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707020032.

Djeddour:2007:TUE

[363] Kh. Djeddour, A. Mokkadem, and M. Pelletier. Test for uniformity by empirical Fourier expansion. Mathematical Methods of Statistics, 16 (2):124–141, June 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530707020044.

Gribkova:2007:EEB

[364] N. V. Gribkova and R. Helmers. On the Edgeworth expansion and the M out of N bootstrap accuracy for a Studentized trimmed mean. Mathematical Methods of Statistics, 16(2):142–176, June 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://

link.springer.com/content/pdf/10.
3103/S1066530707020056.pdf.

Anonymous:2007:HCb

[365] Anonymous. Help & contacts. Mathematical Methods of Statistics, 16(2):??, June 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Arcones:2007:TTM

[366] M. A. Arcones. Two tests for multivariate normality based on the characteristic function. Mathematical Methods of Statistics, 16(3):177–201, September 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707030015.

Guilloux:2007:NEC

[367] A. Guilloux. Nonparametric estimation for censored lifetimes suffering from unknown selection bias. Mathematical Methods of Statistics, 16 (3):202–216, September 2007. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707030027.

Ingster:2007:NHT

[368] Yu. I. Ingster and Yu. A. Kutoyants. Nonparametric hypothesis testing for intensity of the Poisson process. Mathematical Methods of Statistics, 16(3):217–245, September 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707030039.

Lounici:2007:GMA

[369] K. Lounici. Generalized mirror averaging and D-convex aggregation. Mathematical Methods of Statistics, 16 (3):246–259, September 2007. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707030040.

Rigollet:2007:LCA

[370] Ph. Rigollet and A. B. Tsybakov. Linear and convex aggregation of density estimators. Mathematical Methods of Statistics, 16(3):260–280, September 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707030052.

Anonymous:2007:HCc

[371] Anonymous. Help & contacts. Mathematical Methods of Statistics, 16(3): ??, September 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Butler:2007:BAE

[372] L. T. Butler and B. Levit. A Bayesian approach to the estimation of maps between Riemannian manifolds. Mathematical Methods of Statistics, 16(4):281–297, December 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707040011.

Dabo-Niang:2007:KRE

[373] S. Dabo-Niang and A.-F. Yao. Kernel regression estimation for continuous spatial processes. Mathematical Methods of Statistics, 16(4):298–317, December 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707040023.

Ingster:2007:EDH

[374] Yu. Ingster and I. Suslina. Estimation and detection of high-variable functions from Sloan–Woźniakowski space. Mathematical Methods of Statistics, 16(4):318–353, December 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707040035.

Meziani:2007:NEP

[375] K. Méziani. Nonparametric estimation of the purity of a quantum state in quantum homodyne tomography with noisy data. Mathematical Methods of Statistics, 16(4):354–368, December 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707040047.

Aly:2007:NTC

[376] E.-E. Aly. Nonparametric tests for a change in the coefficient of variation. Mathematical Methods of Statistics, 16(4):369–375, December 2007.

CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530707040059.

Anonymous:2007:HCd

[377] Anonymous. Help & contacts. Mathematical Methods of Statistics, 16(4): ??, December 2007. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Belitser:2008:EBT

[378] E. Belitser and F. Enikeeva. Empirical Bayesian test of the smoothness. Mathematical Methods of Statistics, 17(1):1–18, March 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708010018.

Cho:2008:CSN

[379] J. Cho and B. Levit. Cardinal splines in nonparametric regression. Mathematical Methods of Statistics, 17(1):19–34, March 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653070801002X.

Mnatsakanov:2008:NER

[380] R. Mnatsakanov, L. L. Ruymgaart, and F. H. Ruymgaart. Nonparametric estimation of ruin probabilities given a random sample of claims. Mathematical Methods of Statistics, 17(1):35–43, March 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708010031.

Dippon:2008:EES

[381] J. Dippon. Edgeworth expansions for stochastic approximation theory. Mathematical Methods of Statistics, 17(1):44–65, March 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        3103/S1066530708010043.

Bernardoff:2008:GVE

[382] Ph. Bernardoff, C. Kokonendji, and B. Puig. Generalized variance estimators in the multivariate gamma models. Mathematical Methods of Statistics, 17(1):66–73, March 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530708010055.

Sukhanova:2008:TIT

[383] E. M. Sukhanova. A test for independence of two multivariate samples. Mathematical Methods of Statistics, 17(1):74–86, March 2008. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530708010067.

Anonymous:2008:HCa

[384] Anonymous. Help & contacts. Mathematical Methods of Statistics, 17(1):??, March 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Comte:2008:ADD

[385] F. Comte, J. Dedecker, and M. L. Taupin. Adaptive density deconvolution with dependent inputs. Mathematical Methods of Statistics, 17 (2):87–112, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530708020014.

Gine:2008:ASF

[386] E. Giné and R. Nickl. Adaptation on the space of finite signed measures. Mathematical Methods of Statistics, 17(2):113–122, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708020026.

Bouzebda:2008:TIS

[387] S. Bouzebda and A. Keziou. A test of independence in some copula models. Mathematical Methods of Statistics, 17(2):123–137, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708020038.

Dippon:2008:AER

[388] J. Dippon. Asymptotic expansions of the Robbins–Monro process. Mathematical Methods of Statistics, 17 (2):138–145, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S106653070802004X.

Mokkadem:2008:LMD

[389] A. Mokkadem, M. Pelletier, and B. Thiam. Large and moderate deviations principles for kernel estimators of the multivariate regression. Mathematical Methods of Statistics, 17(2):146–172, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530708020051.

Nahtman:2008:SPI

[390] T. Nahtman and D. von Rosen. Shift permutation invariance in linear random factor models. Mathematical Methods of Statistics, 17(2):173–185, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530708020063.

Anonymous:2008:HCb

[391] Anonymous. Help & contacts. Mathematical Methods of Statistics, 17(2):??, June 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Korostelev:2008:MCC

[392] A. Korostelev and O. Lepski. On a multi-channel change-point problem. Mathematical Methods of Statistics, 17(3):187–197, September 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530708030010.

Holdai:2008:IRM

[393] V. Holdai and A. Korostelev. Image reconstruction in multi-channel model under Gaussian noise. Mathematical Methods of Statistics, 17(3):198–208, September 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530708030022.

Shimizu:2008:SSJ

[394] Ya. Shimizu. Statistical specification of jumps under semiparametric semimartingale models. Mathematical Methods of Statistics, 17(3):209–227, September 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708030034.

Xie:2008:GAM

[395] Y. Xie, J. Yu, and B. Ranneby.

A general autoregressive model with Markov switching: Estimation and consistency. Mathematical Methods of Statistics, 17(3):228–240, September 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708030046.

Kiwitt:2008:ELE

[396] S. Kiwitt, E.-R. Nagel, and N. Neumeyer.

Empirical likelihood estimators for the error distribution in nonparametric regression models. Mathematical Methods of Statistics, 17(3):241–260, September 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708030058.

Mnatsakanov:2008:NNE

[397] R. M. Mnatsakanov, N. Misra, Sh.

Li, and E. J. Harner. K_{n}-nearest neighbor estimators of entropy. Mathematical Methods of Statistics, 17 (3):261–277, September 2008. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653070803006X.

Anonymous:2008:HCc

[398] Anonymous. Help & contacts. Mathematical Methods of Statistics, 17(3):

??, September 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Alquier:2008:PBB

[399] P. Alquier. PAC–Bayesian bounds for randomized empirical risk minimizers. Mathematical Methods of Statistics, 17(4):279–304, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708040017.

Bibi:2008:PGP

[400] A. Bibi and A. Aknouche. On periodic GARCH processes: Stationarity, existence of moments and geometric ergodicity. Mathematical Methods of Statistics, 17(4):305–316, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708040029.

Chesneau:2008:STR

[401] Ch. Chesneau and M. Hebiri. Some theoretical results on the grouped variables lasso. Mathematical Methods of Statistics, 17(4):317–326, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708040030.

Durot:2008:MNR

[402] C. Durot. Monotone nonparametric regression with random design. Mathematical Methods of Statistics, 17(4):327–341, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708040042.

deGunst:2008:ABB

[403] M. C. M. de Gunst and O. Shcherbakova.

Asymptotic behavior of Bayes estimators for hidden Markov models with application to ion channels. Mathematical Methods of Statistics, 17(4):342–356, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708040054.

Srivastava:2008:MKP

[404] M. S. Srivastava, T. von Rosen, and D. von Rosen. Models with a Kronecker product covariance structure: Estimation and testing. Mathematical Methods of Statistics, 17(4):357–370, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530708040066. See com-

ment [434] and response [435].

Anonymous:2008:HCd

[405] Anonymous. Help & contacts. Mathematical Methods of Statistics, 17(4): ??, December 2008. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:2009:THD

[406] M. V. Boldin and I. G. Erlikh. Testing hypotheses on the “drift” of parameters in ARMA and ARCH models. Mathematical Methods of Statistics, 18(1):1–20, March 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709010013.

Castillo:2009:EDR

[407] I. Castillo and J.-M. Loubes. Estimation of the distribution of random shifts deformation. Mathematical Methods of Statistics, 18(1):21–42, March 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530709010025.

Fourdrinier:2009:TSE

[408] D. Fourdrinier, V. Konev, and S. Pergamenshchikov. Truncated sequential estimation of the parameter of a first order autoregressive process with dependent noises. Mathematical Methods of Statistics, 18(1):43–58, March 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530709010037.

Lerasle:2009:ADE

[409] M. Lerasle. Adaptive density estimation of stationary β-mixing and τmixing processes. Mathematical Methods of Statistics, 18(1):59–83, March 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530709010049.

Weba:2009:QWL

[410] M. Weba. A quantitative weak law of large numbers and its application to the delta method. Mathematical Methods of Statistics, 18(1):84–95, March 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530709010050.

Anonymous:2009:HCa

[411] Anonymous. Help & contacts. Mathematical Methods of Statistics, 18(1):??, March 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Dette:2009:KTI

[412] H. Dette and B. Hetzler. Khmaladze transformation of integrated variance processes with applications to goodnessof-fit testing. Mathematical Methods of Statistics, 18(2):97–116, June 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653070902001X.

Diop:2009:RGH

[413] A. Diop and G. S. Lô. Ratio of generalized Hill’s estimator and its asymptotic normality theory. Mathematical Methods of Statistics, 18(2):117–133, June 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709020021.

Jordan:2009:ODP

[414] A. Jordan and B. G. Ivanoff. Onedimensional p–p plots and precedence tests for point processes on R^{d}. Mathematical Methods of Statistics, 18(2):134–158, June 2009. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709020033.

Maillot:2009:ULL

[415] B. Maillot and V. Viallon. Uniform limit laws of the logarithm for nonparametric estimators of the regression function in presence of censored data. Mathematical Methods of Statistics, 18(2):159–184, June 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709020045.

Plakhov:2009:SAA

[416] A. Plakhov and P. Cruz. A stochastic approximation algorithm with multiplicative step size modification. Mathematical Methods of Statistics, 18(2):185–200, June 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709020057.

Anonymous:2009:HCb

[417] Anonymous. Help & contacts. Mathematical Methods of Statistics, 18(2):??, June 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Bar-Lev:2009:RPS

[418] S. K. Bar-Lev and A. M. Kagan. Regression of polynomial statistics on the sample mean and natural exponential families. Mathematical Methods of Statistics, 18(3):201–206, September 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709030016.

Butler:2009:BAE

[419] L. T. Butler and B. Levit. A Bayesian approach to the estimation of maps between Riemannian manifolds. II: Examples. Mathematical Methods of Statistics, 18(3):207–230, September 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709030028.

Chebana:2009:PEC

[420] F. Chebana. Parametric estimation with a class of M-estimators. Mathematical Methods of Statistics, 18 (3):231–240, September 2009. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653070903003X.

Ingster:2009:MGF

[421] Yu. I. Ingster and T. Sapatinas. Minimax goodness-of-fit testing in multivariate nonparametric regression. Mathematical Methods of Statistics, 18 (3):241–269, September 2009. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709030041.

Tse:2009:CQR

[422] S. M. Tse. On the cumulative quantile regression process. Mathematical Methods of Statistics, 18(3):270–279, September 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709030053.

Anonymous:2009:HCc

[423] Anonymous. Help & contacts. Mathematical Methods of Statistics, 18(3): ??, September 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Buchmann:2009:WEP

[424] B. Buchmann. Weighted empirical processes in the nonparametric inference for Lévy processes. Mathematical Methods of Statistics, 18(4):281–309, December 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709040012.

Ingster:2009:EDF

[425] Yu. I. Ingster and N. Stepanova. Estimation and detection of functions from weighted tensor product spaces. Mathematical Methods of Statistics, 18(4):310–340, December 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709040024.

Plancade:2009:EDR

[426] S. Plancade. Estimation of the density of regression errors by pointwise model selection. Mathematical Methods of Statistics, 18(4):341–374, December 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709040036.

Rukhin:2009:CCE

[427] A. L. Rukhin. Conservative confidence ellipsoids for linear model parameters. Mathematical Methods of Statistics, 18(4):375–396, December 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530709040048.

Anonymous:2009:HCd

[428] Anonymous. Help & contacts. Mathematical Methods of Statistics, 18(4): ??, December 2009. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Abdi:2010:CNC

[429] S. A. Ould Abdi, S. Dabo-Niang, A. Diop, and A. Ould Abdi. Consistency of a nonparametric conditional quantile estimator for random fields. Mathematical Methods of Statistics, 19(1):1–21, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710010011.

Bordes:2010:STC

[430] L. Bordes and P. Vandekerkhove.

Semiparametric two-component mixture model with a known component: An asymptotically normal estimator. Mathematical Methods of Statistics, 19(1):22–41, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710010023.

Comte:2010:AEC

[431] F. Comte and J. Johannes. Adaptive estimation in circular functional linear models. Mathematical Methods of Statistics, 19(1):42–63, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710010035.

Meintanis:2010:TSN

[432] S. G. Meintanis. Testing skew normality via the moment generating function. Mathematical Methods of Statistics, 19(1):64–72, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710010047.

Menendez:2010:PTE

[433] M. L. Menéndez, L. Pardo, and K. Zografos. Preliminary test estimators in intraclass correlation model under unequal family sizes. Mathematical Methods of Statistics, 19(1):73–87, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710010059. See cor-

rection [442].

Lee:2010:CMK

[434] C. H. Lee, P. Dutilleul, and A. Roy. Comment on “Models with a Kronecker Product Covariance Structure: Estimation and Testing” by M. S. Srivastava, T. von Rosen, and D. von Rosen, Mathematical Methods of Statistics, 17 (2008), pp. 357–370. Mathematical Methods of Statistics, 19(1):88–90, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710010060.  See

[404].

Srivastava:2010:RCC

[435] M. S. Srivastava, T. von Rosen, and D. von Rosen. Response to comment by C. H. Lee, P. Dutilleul and A. Roy on “Models with a Kronecker Product Covariance Structure: Estimation and Testing”. Mathematical Methods of Statistics, 19(1):91, March 2010. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        3103/S1066530710010072.pdf.  See

[404, 434].

Anonymous:2010:HCa

[436] Anonymous. Help & contacts. Mathematical Methods of Statistics, 19(1):??, March 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Balan:2010:AOE

[437] R. M. Balan, L. Dumitrescu, and I. Schiopu-Kratina. Asymptotically optimal estimating equation with strongly consistent solutions for longitudinal data. Mathematical Methods of Statistics, 19(2):93–120, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710020018.

Burke:2010:AMH

[438] M. D. Burke. Approximations for a multivariate hybrid process with applications to change-point detection. Mathematical Methods of Statistics, 19(2):121–135, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071002002X.

Laloe:2010:QCB

[439] T. Laloë. L_{1}-quantization and clustering in Banach spaces. Mathematical Methods of Statistics, 19(2):136–150, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710020031.

DiNardo:2010:NAS

[440] E. Di Nardo. A new approach to Sheppard’s corrections. Mathematical Methods of Statistics, 19(2):151–162, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710020043.

Sharia:2010:ELE

[441] T. Sharia. Efficient on-line estimation of autoregressive parameters. Mathematical Methods of Statistics, 19(2):163–186, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710020055.

Menendez:2010:CPT

[442] M. L. Menéndez, L. Pardo, and K. Zografos. Correction to: “Preliminary test estimators in intraclass correlation model under unequal family sizes”. Mathematical Methods of Statistics, 19(2):187, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/content/pdf/10.3103/
                                        S1066530710020067.pdf. See [433].

Anonymous:2010:HCb

[443] Anonymous. Help & contacts. Mathematical Methods of Statistics, 19(2):??, June 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Akakpo:2010:HSP

[444] N. Akakpo and C. Durot. Histogram selection for possibly censored data. Mathematical Methods of Statistics, 19(3):189–218, September 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710030014.

Babenko:2010:OCR

[445] A. Babenko and E. Belitser. Oracle convergence rate of posterior under projection prior and Bayesian model selection. Mathematical Methods of Statistics, 19(3):219–245, September 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710030026.

Ferger:2010:MDE

[446] D. Ferger. Minimum distance estimation in normed linear spaces with donsker-classes. Mathematical Methods of Statistics, 19(3):246–266, September 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710030038.

Kaehler:2010:GSS

[447] B. D. Kaehler and R. A. Maller.

A generalized skewness statistic for stationary ergodic martingale differences. Mathematical Methods of Statistics, 19(3):267–282, September 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071003004X.

Levit:2010:MR

[448] B. Levit. Minimax revisited. I. Mathematical Methods of Statistics, 19 (3):283–297, September 2010. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710030051.

Anonymous:2010:HCc

[449] Anonymous. Help & contacts. Mathematical Methods of Statistics, 19(3): ??, September 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Levit:2010:MRI

[450] B. Levit. Minimax revisited. II. Mathematical Methods of Statistics, 19 (4):299–326, December 2010. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710040010.

Nickl:2010:ESB

[451] R. Nickl and B. M. Pötscher. Efficient simulation-based minimum distance estimation and indirect inference. Mathematical Methods of Statistics, 19(4):327–364, December 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710040022.

Holdai:2010:DSE

[452] V. Holdai and A. Korostelev. Detection of slightly expressed changes in random environment. Mathematical Methods of Statistics, 19(4):365–373, December 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530710040034.

Dutta:2010:MED

[453] S. Dutta and A. Goswami. Mode estimation for discrete distributions. Mathematical Methods of Statistics, 19 (4):374–384, December 2010. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530710040046.

Anonymous:2010:HCd

[454] Anonymous. Help & contacts. Mathematical Methods of Statistics, 19(4): ??, December 2010. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Boldin:2011:LRS

[455] M. V. Boldin. Local robustness of sign tests in AR (1) against outliers. Mathematical Methods of Statistics, 20(1):1–13, March 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711010017.

Bouzebda:2011:SPU

[456] S. Bouzebda, A. Keziou, and T. Zari.

K-sample problem using strong approximations of empirical copula processes. Mathematical Methods of Statistics, 20(1):14–29, March 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711010029.

Guillou:2011:MPE

[457] A. Guillou and N. Klutchnikoff. Minimax pointwise estimation of an anisotropic regression function with unknown density of the design. Mathematical Methods of Statistics, 20 (1):30–57, March 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530711010030.

Hopfner:2011:EPP

[458] R. Höpfner and Yu. Kutoyants. Estimating a periodicity parameter in the drift of a time inhomogeneous diffusion. Mathematical Methods of Statistics, 20(1):58–74, March 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530711010042.

Kashitsyn:2011:MMK

[459] P. A. Kashitsyn. Multivariate model with a Kronecker product covariance structure: S. N. Roy method of estimation. Mathematical Methods of Statistics, 20(1):75–78, March 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530711010054.

Anonymous:2011:HCa

[460] Anonymous. Help & contacts. Mathematical Methods of Statistics, 20(1):??, March 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Bertail:2011:RAM

[461] P. Bertail and S. Clémençon. A renewal approach to Markovian Ustatistics. Mathematical Methods of Statistics, 20(2):79–105, June 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530711020013.

Lejeune:2011:PLD

[462] F.-X. Lejeune. Piecewise linear density estimation for sampled data. Mathematical Methods of Statistics, 20 (2):106–124, June 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530711020025.

Shimizu:2011:EED

[463] Y. Shimizu. Estimation of the expected discounted penalty function for Lévy insurance risks. Mathematical Methods of Statistics, 20(2):125–149, June 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711020037.

Khan:2011:EIL

[464] R. U. Khan and D. Kumar. Expectation identities of lower generalized order statistics from generalized exponential distribution and a characterization. Mathematical Methods of Statistics, 20(2):150–157, June 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711020049.

Tahata:2011:LOQ

[465] K. Tahata, H. Yamamoto, and S. Tomizawa. Linear ordinal quasisymmetry model and decomposition of symmetry for multi-way tables. Mathematical Methods of Statistics, 20(2):158–164, June 2011. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711020050.

Wang:2011:ELP

[466] L. Wang. L_{1}-estimation for the location parameters in stochastic volatility models. Mathematical Methods of Statistics, 20(2):165–170, June 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530711020062.

Anonymous:2011:HCb

[467] Anonymous. Help & contacts. Mathematical Methods of Statistics, 20(2):??, June 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Bobkov:2011:NUBa

[468] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Non-uniform bounds in local limit theorems in case of fractional moments. I. Mathematical Methods of Statistics, 20(3):171–191, September 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S106653071103001X.

Bouzebda:2011:SNM

[469] S. Bouzebda. Some new multivariate tests of independence. Mathematical Methods of Statistics, 20 (3):192–205, September 2011. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530711030021.

Espinasse:2011:EEB

[470] T. Espinasse, F. Gamboa, and J.- M. Loubes. Estimation error for blind Gaussian time series prediction. Mathematical Methods of Statistics, 20(3):206–223, September 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530711030033.

Hirose:2011:ALE

[471] Y. Hirose. Asymptotic linear expansion of profile likelihood in the Cox mode. Mathematical Methods of Statistics, 20(3):224–231, September 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711030045.

Gaines:2011:ADM

[472] G. Gaines, K. Kaphle, and F. Ruymgaart. Application of a deltamethod for random operators to testing equality of two covariance operators. Mathematical Methods of Statistics, 20(3):232–245, September 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711030057.

Yode:2011:AMT

[473] A. F. Yodé. Adaptive minimax test of independence. Mathematical Methods of Statistics, 20(3):246–268, September 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711030069.

Anonymous:2011:HCc

[474] Anonymous. Help & contacts. Mathematical Methods of Statistics, 20(3): ??, September 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Bobkov:2011:NUBb

[475] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Non-uniform bounds in local limit theorems in case of fractional moments. II. Mathematical Methods of Statistics, 20(4):269–287, December 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711040016.

Gach:2011:NML

[476] F. Gach and B. M. Pötscher. Nonparametric maximum likelihood density estimation and simulation-based minimum distance estimators. Mathematical Methods of Statistics, 20 (4):288–326, December 2011. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711040028.

Gronneberg:2011:ECS

[477] S. Grønneberg and N. L. Hjort. On the errors committed by sequences of estimator functionals. Mathematical Methods of Statistics, 20(4):327–346, December 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071104003X.

Ingster:2011:MNT

[478] Yu. I. Ingster, T. Sapatinas, and I. A. Suslina. Minimax nonparametric testing in a problem related to the Radon transform. Mathematical Methods of Statistics, 20(4):347–364, December 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711040041.

Barakat:2011:LDI

[479] H. M. Barakat and A. R. Omar. On limit distributions for intermediate order statistics under power normalization. Mathematical Methods of Statistics, 20(4):365–377, December 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530711040053.

Anonymous:2011:HCd

[480] Anonymous. Help & contacts. Mathematical Methods of Statistics, 20(4): ??, December 2011. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Akakpo:2012:AAI

[481] N. Akakpo. Adaptation to anisotropy and inhomogeneity via dyadic piecewise polynomial selection. Mathematical Methods of Statistics, 21(1):1–28, March 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712010012.

Enikeeva:2012:TER

[482] F. Enikeeva. On two estimates related to the change-point problem. Mathematical Methods of Statistics, 21(1):29–42, March 2012. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712010024.

Kharin:2012:APP

[483] Yu. S. Kharin and V. A. Voloshko.

On asymptotic properties of the plugin cepstrum estimator for Gaussian time series. Mathematical Methods of Statistics, 21(1):43–60, March 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712010036.

Ushakov:2012:NSD

[484] N. G. Ushakov. A note on superkernel density estimators. Mathematical Methods of Statistics, 21(1):61–68, March 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530712010048.

Anonymous:2012:HCa

[485] Anonymous. Help & contacts. Mathematical Methods of Statistics, 21(1):??, March 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Carabarin-Aguirre:2012:PDE

[486] A. Carabarin-Aguirre and B. G. Ivanoff. Path-dependent estimation of a distribution under generalized censoring. Mathematical Methods of Statistics, 21(2):69–92, April 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530712020019.

Autin:2012:ATC

[487] F. Autin and C. Pouet. Adaptive test on components of densities mixture. Mathematical Methods of Statistics, 21(2):93–108, April 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530712020020.

Wagener:2012:BEA

[488] J. Wagener and H. Dette. Bridge estimators and the adaptive lasso under heteroscedasticity. Mathematical Methods of Statistics, 21(2):109–126, April 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712020032.

Wagener:2012:QPU

[489] J. Wagener, S. Volgushev, and H. Dette.

The quantile process under random censoring. Mathematical Methods of Statistics, 21(2):127–141, April 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712020044.

Chigansky:2012:NTP

[490] P. Chigansky and Yu. A. Kutoyants. On nonlinear TAR processes and threshold estimation. Mathematical Methods of Statistics, 21(2):142–152, April 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712020056.

Anonymous:2012:HCb

[491] Anonymous. Help & contacts. Mathematical Methods of Statistics, 21(2):??, April 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Bouzebda:2012:SAB

[492] S. Bouzebda. On the strong approximation of bootstrapped empirical copula processes with applications. Mathematical Methods of Statistics, 21(3):153–188, July 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712030015.

Johannes:2012:AEL

[493] J. Johannes and R. Schenk. Adaptive estimation of linear functionals in functional linear models. Mathematical Methods of Statistics, 21 (3):189–214, July 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530712030027.

Niang:2012:WBE

[494] M. A. Niang, G. M. Nkiet, and A. Diop. Wavelets-based estimation of nonlinear canonical analysis. Mathematical Methods of Statistics, 21(3):215–237, July 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712030039.

Tse:2012:CQR

[495] S. M. Tse. The cumulative quantile regression function with censored and truncated covariate. Mathematical Methods of Statistics, 21 (3):238–249, July 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530712030040.

Anonymous:2012:HCc

[496] Anonymous. Help & contacts. Mathematical Methods of Statistics, 21(3):??, July 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Barnard:2012:BCS

[497] R. W. Barnard and P. Hadjicostas. Banks’ criterion and symmetric stable laws with index of stability between one-half and one. Mathematical Methods of Statistics, 21(4):251–282, October 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712040011.

Biscay:2012:ACE

[498] R. Biscay, H. Lescornel, and J.-M.

Loubes. Adaptive covariance estimation with model selection. Mathematical Methods of Statistics, 21 (4):283–297, October 2012. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712040023.

Louani:2012:LDR

[499] D. Louani and S. M. Ould Maouloud.

Large deviation results for the nonparametric regression function estimator on functional data. Mathematical Methods of Statistics, 21(4):298–313, October 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530712040035.

Anonymous:2012:HCd

[500] Anonymous. Help & contacts. Mathematical Methods of Statistics, 21(4):??, October 2012. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Lepski:2013:UFPa

[501] O. Lepski. Upper functions for positive random functionals. I. General setting and Gaussian random functions. Mathematical Methods of Statistics, 22(1):1–27, January 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530713010018.

Duval:2013:NER

[502] C. Duval. Nonparametric estimation of a renewal reward process from discrete data. Mathematical Methods of Statistics, 22(1):28–56, January 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S106653071301002X.

Agostinelli:2013:ASP

[503] C. Agostinelli and M. Romanazzi. Asymptotics of stationary points of local simplicial depth in the univariate case. Mathematical Methods of Statistics, 22(1):57–69, January 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530713010031.

Nkurunziza:2013:BRF

[504] S. Nkurunziza. The bias and risk functions of some Stein-rules in elliptically contoured distributions. Mathematical Methods of Statistics, 22 (1):70–82, January 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        3103/S1066530713010043.

Anonymous:2013:HCa

[505] Anonymous. Help & contacts. Mathematical Methods of Statistics, 22(1):??, January 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Lepski:2013:UFPb

[506] O. Lepski. Upper functions for positive random functionals. II. Application to the empirical processes theory, Part 1. Mathematical Methods of Statistics, 22(2):83–99, April 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530713020014.

Spokoiny:2013:SDB

[507] V. Spokoiny and M. Zhilova. Sharp deviation bounds for quadratic forms. Mathematical Methods of Statistics, 22(2):100–113, April 2013. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530713020026.

Stepanova:2013:EAD

[508] N. Stepanova. On estimation of analytic density functions in L_{p}. Mathematical Methods of Statistics, 22 (2):114–136, April 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530713020038.

Wagener:2013:ALH

[509] J. Wagener and H. Dette. The adaptive lasso in high-dimensional sparse heteroscedastic models. Mathematical Methods of Statistics, 22 (2):137–154, April 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S106653071302004X.

Joutard:2013:SLD

[510] C. Joutard. A strong large deviation theorem. Mathematical Methods of Statistics, 22(2):155–164, April 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530713020051.

Anonymous:2013:HCb

[511] Anonymous. Help & contacts. Mathematical Methods of Statistics, 22(2):??, April 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Levit:2013:SNP

[512] B. Levit. Some new perspectives in best approximation and interpolation of random data. Mathematical Methods of Statistics, 22(3):165–192, July 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530713030010.

Lepski:2013:UFPc

[513] O. Lepski. Upper functions for positive random functionals. II. Application to the empirical processes theory, Part 2. Mathematical Methods of Statistics, 22(3):193–212, July 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530713030022.

Cadre:2013:SCD

[514] B. Cadre. Supervised classification of diffusion paths. Mathematical Methods of Statistics, 22(3):213–225, July 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530713030034.

Bouzebda:2013:ABW

[515] S. Bouzebda and T. Zari. Asymptotic behavior of weighted multivariate Cramér–von Mises-type statistics under contiguous alternatives. Mathematical Methods of Statistics, 22 (3):226–252, July 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530713030046.

Anonymous:2013:HCc

[516] Anonymous. Help & contacts. Mathematical Methods of Statistics, 22(3):??, July 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Chagny:2013:WBC

[517] G. Chagny. Warped bases for conditional density estimation. Mathematical Methods of Statistics, 22 (4):253–282, October 2013. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530713040017.

Fasen:2013:SIS

[518] V. Fasen. Statistical inference of spectral estimation for continuoustime MA processes with finite second moments. Mathematical Methods of Statistics, 22(4):283–309, October 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530713040029.

Lepski:2013:AES

[519] O. Lepski and N. Serdyukova. Adaptive estimation in the single-index model via oracle approach. Mathematical Methods of Statistics, 22(4):310–332, October 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530713040030.

Esaulov:2013:REP

[520] D. M. Esaulov. Residual empirical processes and their application to GM-testing for the autoregression order. Mathematical Methods of Statistics, 22(4):333–349, October 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530713040042.

Anonymous:2013:HCd

[521] Anonymous. Help & contacts. Mathematical Methods of Statistics, 22(4):??, October 2013. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Falconnet:2014:ANE

[522] M. Falconnet, D. Loukianova, and C. Matias. Asymptotic normality and efficiency of the maximum likelihood estimator for the parameter of a ballistic random walk in a random environment. Mathematical Methods of Statistics, 23(1):1–19, January 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714010013.

Golubev:2014:CIE

[523] Yu. Golubev and D. Ostrovski. Concentration inequalities for the exponential weighting method. Mathematical Methods of Statistics, 23(1):20–37, January 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714010025.

Bouzebda:2014:APP

[524] S. Bouzebda. Asymptotic properties of pseudo maximum likelihood estimators and test in semi-parametric copula models with multiple change points. Mathematical Methods of Statistics, 23(1):38–65, January 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530714010037.

Tripathi:2014:EPE

[525] Y. M. Tripathi, S. Kumar, and C. Petropoulos. Estimation of the parameters of an exponential distribution under constrained location. Mathematical Methods of Statistics, 23(1):66–79, January 2014. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530714010049.

Anonymous:2014:HCa

[526] Anonymous. Help & contacts. Mathematical Methods of Statistics, 23(1):??, January 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Aknouche:2014:ESS

[527] A. Aknouche. Estimation and strict stationarity testing of ARCH processes based on weighted least squares. Mathematical Methods of Statistics, 23(2):81–102, April 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071402001X.

Brouste:2014:APM

[528] A. Brouste, C. Cai, and M. Kleptsyna. Asymptotic properties of the MLE for the autoregressive process coefficients under stationary Gaussian noise. Mathematical Methods of Statistics, 23(2):103–115, April 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714020021.

Chernousova:2014:SCR

[529] E. Chernousova and Yu. Golubev. Spectral cut-off regularizations for ill-posed linear models. Mathematical Methods of Statistics, 23(2):116–131, April 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714020033.

Klutchnikoff:2014:PAE

[530] N. Klutchnikoff. Pointwise adaptive estimation of a multivariate function. Mathematical Methods of Statistics, 23(2):132–150, April 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714020045.

Tze:2014:CCT

[531] S. M. Tze. Concentration curve for truncated and censored data. Mathematical Methods of Statistics, 23 (2):151–157, April 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        3103/S1066530714020057.

Anonymous:2014:HCb

[532] Anonymous. Help & contacts. Mathematical Methods of Statistics, 23(2):??, April 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Falconnet:2014:MLE

[533] M. Falconnet, A. Gloter, and D. Loukianova.

Maximum likelihood estimation in the context of a sub-ballistic random walk in a parametric random environment. Mathematical Methods of Statistics, 23(3):159–175, July 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530714030016.

Marteau:2014:GRS

[534] C. Marteau and P. Mathé. General regularization schemes for signal detection in inverse problems. Mathematical Methods of Statistics, 23(3):176–200, July 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530714030028.

Muller:2014:EDU

[535] U. Müller and W. Wefelmeyer. Estimating a density under pointwise constraints on the derivatives. Mathematical Methods of Statistics, 23(3):201–209, July 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071403003X.

Bouzebda:2014:SAM

[536] S. Bouzebda and T. Zari. Strong approximation of multidimensional P – P plots processes by Gaussian processes with applications to statistical tests. Mathematical Methods of Statistics, 23(3):210–238, July 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530714030041.

Anonymous:2014:HCc

[537] Anonymous. Help & contacts. Mathematical Methods of Statistics, 23(3):??, July 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Blanke:2014:EBI

[538] D. Blanke and D. Bosq. Exponential bounds for intensity of jumps. Mathematical Methods of Statistics, 23(4):239–255, October 2014. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714040012.

Csorgo:2014:ALB

[539] M. Csörgő, Yu. V. Martsynyuk, and M. M. Nasari. Another look at bootstrapping the Student t-statistic. Mathematical Methods of Statistics, 23(4):256–278, October 2014. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714040024.

Ingster:2014:SDI

[540] Yu. Ingster, B. Laurent, and C. Marteau. Signal detection for inverse problems in a multidimensional framework. Mathematical Methods of Statistics, 23(4):279–305, October 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530714040036.

Slaoui:2014:SAM

[541] Y. Slaoui. The stochastic approximation method for estimation of a distribution function. Mathematical Methods of Statistics, 23(4):306–325, October 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530714040048.

Anonymous:2014:HCd

[542] Anonymous. Help & contacts. Mathematical Methods of Statistics, 23(4):??, October 2014. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Golubev:2015:EIP

[543] Yu. Golubev and T. Zimolo. Estimation in ill-posed linear models with nuisance design. Mathematical Methods of Statistics, 24(1):1–15, January 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715010019.

Hao:2015:EIA

[544] C. Hao, D. von Rosen, and T. von Rosen. Explicit influence analysis in two-treatment balanced crossover models. Mathematical Methods of Statistics, 24(1):16–36, January 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715010020.

Levit:2015:VOD

[545] B. Levit. Variance-optimal data approximation using the Abel–Jacobi functions. Mathematical Methods of Statistics, 24(1):37–54, January 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715010032.

Vancak:2015:CSM

[546] V. Vancak, Y. Goldberg, S. K. Bar-Lev, and B. Boukai. Continuous statistical models: With or without truncation parameters? Mathematical Methods of Statistics, 24(1):55–73, January 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715010044.

Pinelis:2015:MTM

[547] I. Pinelis. Monotone tail and moment ratio properties of Student’s family of distributions. Mathematical Methods of Statistics, 24(1):74–79, January 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715010056.

Anonymous:2015:HC

[548] Anonymous. Help & contacts. Mathematical Methods of Statistics, 24(1):??, January 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Mabon:2015:AEM

[549] G. Mabon. Adaptive estimation of marginal random-effects densities in linear mixed-effects models. Mathematical Methods of Statistics, 24 (2):81–109, April 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        3103/S1066530715020015.

Gushchin:2015:CMT

[550] A. Gushchin. A characterization of maximin tests for two composite hypotheses. Mathematical Methods of Statistics, 24(2):110–121, April 2015.

CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715020027.

Jana:2015:EOS

[551] N. Jana and S. Kumar. Estimation of ordered scale parameters of two exponential distributions with a common guarantee time. Mathematical Methods of Statistics, 24(2):122–134, April 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715020039.

Meyer:2015:BIT

[552] M. Meyer, T. McMurry, and D. Politis. Baxter’s inequality for triangular arrays. Mathematical Methods of Statistics, 24(2):135–146, April 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715020040.

Tse:2015:CQR

[553] S. M. Tse. The cumulative quantile regression function with censored and truncated response. Mathematical Methods of Statistics, 24(2):147–155, April 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715020052.

Wang:2015:CRS

[554] J. Wang. A comprehensive result on stochastic comparison for parallel systems in terms of hazard rate order. Mathematical Methods of Statistics, 24(2):156–161, April 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715020064.

Bouzebda:2015:MWD

[555] S. Bouzebda, S. Didi, and L. El Hajj. Multivariate wavelet density and regression estimators for stationary and ergodic continuous time processes: Asymptotic results. Mathematical Methods of Statistics, 24(3):163–199, July 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715030011.

Rebafka:2015:NEM

[556] T. Rebafka and F. Roueff. Nonparametric estimation of the mixing density using polynomials. Mathematical Methods of Statistics, 24(3):200–224, July 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715030023.

Kumar:2015:EES

[557] D. Kumar. Explicit expressions and statistical inference of generalized Rayleigh distribution based on lower record values. Mathematical Methods of Statistics, 24(3):225–241, July 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715030035.

Belitser:2015:RTA

[558] E. Belitser and P. Serra. Recursive tracking algorithm for a predictable time-varying parameter of a time series. Mathematical Methods of Statistics, 24(4):243–265, October 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715040018.

Brahimi:2015:GAE

[559] B. Brahimi, D. Meraghni, and A. Necir.

Gaussian approximation to the extreme value index estimator of a heavy-tailed distribution under random censoring. Mathematical Methods of Statistics, 24(4):266–279, October 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071504002X.

Nzabanita:2015:EGM

[560] J. Nzabanita, D. von Rosen, and M. Singull. Extended GMANOVA model with a linearly structured covariance matrix. Mathematical Methods of Statistics, 24(4):280–291, October 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715040031.

Rukhin:2015:HOA

[561] A. Rukhin. Higher order accurate procedures to compare two normal populations. Mathematical Methods of Statistics, 24(4):292–308, October 2015. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530715040043.

Yaskov:2015:VIQ

[562] P. Yaskov. Variance inequalities for quadratic forms with applications. Mathematical Methods of Statistics, 24(4):309–319, October 2015.

CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530715040055.

Golubev:2016:TPR

[563] Yu. Golubev and Th. Zimolo. Tikhonov– Phillips regularizations in linear models with blurred design. Mathematical Methods of Statistics, 25 (1):1–25, January 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        3103/S1066530716010014.

Rebelles:2016:SAD

[564] G. Rebelles. Structural adaptive deconvolution under L_{p}-losses. Mathematical Methods of Statistics, 25 (1):26–53, January 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        3103/S1066530716010026.

Nikitin:2016:EET

[565] Ya. Yu. Nikitin and K. Yu. Volkova. Efficiency of exponentiality tests based on a special property of exponential distribution. Mathematical Methods of Statistics, 25(1):54–66, January 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530716010038.

Savelov:2016:EPH

[566] M. P. Savelov. Extremal problems for hypotheses testing with set-valued decisions. Mathematical Methods of Statistics, 25(1):67–77, January 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071601004X.

Bar-Lev:2016:OBI

[567] S. K. Bar-Lev, D. Bshouty, and F. A.

van der Duyn Schouten. Operatorbased intensity functions for the nonhomogeneous Poisson process. Mathematical Methods of Statistics, 25(2): 79–98, April 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716020010.

Kutoyants:2016:SFG

[568] Yu. A. Kutoyants. On score-functions and goodness-of-fit tests for stochastic processes. Mathematical Methods of Statistics, 25(2):99–120, April 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.3103/
S1066530716020022.

Ahmad:2016:TSI

[569] M. R. Ahmad. On testing sphericity and identity of a covariance matrix with large dimensions. Mathematical Methods of Statistics, 25 (2):121–132, April 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716020034.

Groenitz:2016:ROE

[570] H. Groenitz. Risk-optimal estimators for survey procedures with certain indirect questions. Mathematical Methods of Statistics, 25(2):133– 144, April 2016. CODEN ????

ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716020046.

Wang:2016:LRO

[571] J. Wang and P. Zhao. On likelihood ratio ordering of parallel systems with exponential components. Mathematical Methods of Statistics, 25 (2):145–150, April 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716020058.

Mokkadem:2016:MRO

[572] A. Mokkadem and M. Pelletier. The multivariate Révész’s online estimator of a regression function and its averaging. Mathematical Methods of Statistics, 25(3):151–167, July 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716030017.

Bouzebda:2016:LLR

[573] S. Bouzebda, M. Chaouch, and N. Laïb. Limiting law results for a class of conditional mode estimates for functional stationary ergodic data. Mathematical Methods of Statistics, 25 (3):168–195, July 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716030029.

Hung:2016:CSE

[574] Y. C. Hung, R. W. Chen, and N. Balakrishnan. On the correlation structure of exponential order statistics and some extensions. Mathematical Methods of Statistics, 25 (3):196–206, July 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716030030.

Kleptsyna:2016:RDT

[575] M. L. Kleptsyna and A. Yu. Veretennikov. On robustness of discrete time optimal filters. Mathematical Methods of Statistics, 25(3):207– 218, July 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716030042.

Martsynyuk:2016:TCM

[576] Yu. V. Martsynyuk. Testing for change in the mean via convergence in distribution of sup-functionals of weighted tied-down partial sums processes. Mathematical Methods of Statistics, 25(3):219–232, July 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716030054.

Haywood:2016:BCE

[577] J. Haywood. A brief comment on exponentiality and reign lengths of emperors. Mathematical Methods of Statistics, 25(3):233–234, July 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716030066.

Levit:2016:OMI

[578] B. Levit. Optimal methods of interpolation in nonparametric regression.

Mathematical Methods of Statistics, 25 (4):235–261, October 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716040013.

Sharia:2016:RCT

[579] T. Sharia and L. Zhong. Rate of convergence of truncated stochastic approximation procedures with moving bounds. Mathematical Methods of Statistics, 25 (4):262–280, October 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716040025.

Bouzebda:2016:SAS

[580] S. Bouzebda. Some applications of the strong approximation of the integrated empirical copula processes. Mathematical Methods of Statistics, 25 (4):281–303, October 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716040037.

Gushchin:2016:MID

[581] A. A. Gushchin. The minimum increment of f-divergences given total variation distances. Mathematical Methods of Statistics, 25(4):304– 312, October 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.3103/S1066530716040049.

Gribkova:2016:CTM

[582] N. Gribkova. Cramér type moderate deviations for trimmed L-statistics.

Mathematical Methods of Statistics, 25 (4):313–322, October 2016. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http://link.

springer.com/accesspage/article/
10.3103/S1066530716040050.

Gasparyan:2017:LBS

[583] S. B. Gasparyan and Y. A. Kutoyants.

On the lower bound in second order estimation for Poisson processes: Asymptotic efficiency. Mathematical Methods of Statistics, 26(1):1–19, January 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Gatto:2017:SAD

[584] R. Gatto. Saddlepoint approximations to the distribution of the total distance of the multivariate isotropic and von Mises–Fisher random walks. Mathematical Methods of Statistics, 26(1):20–36, January 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Sharia:2017:ABT

[585] T. Sharia and L. Zhong. Asymptotic behavior of truncated stochastic approximation procedures. Mathematical Methods of Statistics, 26(1):37–54, January 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Alquier:2017:OIQ

[586] P. Alquier and B. Guedj. An oracle inequality for quasi-Bayesian nonnegative matrix factorization. Mathematical Methods of Statistics, 26(1):55–67, January 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Fall:2017:SAC

[587] A. M. Fall, G. S. Lo, A. Adekpedjou, and C. H. Ndiaye. A supermartingale argument for characterizing the functional Hill process weak law for small parameters. Mathematical Methods of Statistics, 26(1):68–80, January 2017. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Masuda:2017:MCR

[588] H. Masuda and Y. Shimizu. Moment convergence in regularized estimation under multiple and mixed-rates asymptotics. Mathematical Methods of Statistics, 26(2):81–110, April 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Dedecker:2017:MWS

[589] J. Dedecker and G. Saulière. The Mann– Whitney U-statistic for α-dependent sequences. Mathematical Methods of Statistics, 26(2):111–133, April 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Kubokawa:2017:UAE

[590] T. Kubokawa, É. Marchand, and W. E. Strawderman. A unified approach to estimation of noncentrality parameters, the multiple correlation coefficient, and mixture models. Mathematical Methods of Statistics, 26(2):134–148, April 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Mori:2017:RUR

[591] T. F. Móri and G.-J. Székely. Representations by uncorrelated random variables. Mathematical Methods of Statistics, 26(2):149–153, April 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Zinger:2017:UCS

[592] A. A. Zinger. On unlinking of continuous statistics of normal sample. Mathematical Methods of Statistics, 26(2):154–158, April 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Bar-Lev:2017:MVP

[593] S. K. Bar-Lev and C. C. Kokonendji. On the mean value parametrization of natural exponential families — a revisited review. Mathematical Methods of Statistics, 26(3):159–175, July 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Schick:2017:EEE

[594] A. Schick and Y. Zhu. Efficient estimation of the error distribution in a varying coefficient regression model. Mathematical Methods of Statistics, 26(3):176–195, July 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Nkiet:2017:ATM

[595] G. M. Nkiet. Asymptotic theory of multiple-set linear canonical analysis. Mathematical Methods of Statistics, 26 (3):196–211, July 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Al-Labadi:2017:TSK

[596] L. Al-Labadi and M. Zarepour. Twosample Kolmogorov–Smirnov test using a Bayesian nonparametric approach. Mathematical Methods of Statistics, 26 (3):212–225, July 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Patra:2017:CIE

[597] L. K. Patra and S. Kumar. Classes of improved estimators for parameters of a Pareto distribution. Mathematical Methods of Statistics, 26(3):226–235, July 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Comte:2017:LDU

[598] F. Comte and G. Mabon. Laguerre deconvolution with unknown matrix operator. Mathematical Methods of Statistics, 26(4):237–266, October 2017. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Gribkova:2017:SFA

[599] N. Gribkova and R. Zitikis. Statistical foundations for assessing the difference between the classical and weighted-Gini betas. Mathematical Methods of Statistics, 26(4):267–281, October 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Marteau:2017:MSD

[600] C. Marteau and Th. Sapatinas. Minimax signal detection under weak noise assumptions. Mathematical Methods of Statistics, 26(4):282–298, October 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

vonRosen:2017:ESR

[601] T. von Rosen and D. von Rosen. On estimation in some reduced rank extended growth curve models. Mathematical Methods of Statistics, 26(4): 299–310, October 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Balakrishnan:2017:JWR

[602] N. Balakrishnan, G. Barmalzan, and S. Kosari. On joint weak reversed hazard rate order under symmetric copulas. Mathematical Methods of Statistics, 26(4):311–318, October 2017. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic).

Maillard:2018:BCP

[603] O.-A. Maillard. Boundary crossing probabilities for general exponential families. Mathematical Methods of Statistics, 27(1):1–31, January 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718010015.

Nagler:2018:AAJ

[604] T. Nagler. Asymptotic analysis of the jittering kernel density estimator. Mathematical Methods of Statistics, 27(1):32–46, January 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718010027.

Barakat:2018:LTB

[605] H. M. Barakat, M. A. Abd Elgawad, H. Qin, and T. Yan. Limit theory of bivariate generalized order statistics with random sample size. Mathematical Methods of Statistics, 27(1):47–59, January 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718010039.

Molchanov:2018:EMD

[606] S. Molchanov, Z. Zhang, and L. Zheng. Entropic moments and domains of attraction on countable alphabets. Mathematical Methods of Statistics, 27(1):60–70, January 2018. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718010040.

Tsai:2018:MLE

[607] Ming-Tien Tsai. On the maximum likelihood estimation of a covariance matrix. Mathematical Methods of Statistics, 27(1):71–82, January 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718010052.

Chen:2018:EII

[608] L. Chen, Y. Davydov, N. Gribkova, and R. Zitikis. Estimating the index of increase via balancing deterministic and random data. Mathematical Methods of Statistics, 27(2):83–102, April 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718020011.

Kharin:2018:SEP

[609] Yu. S. Kharin, V. A. Voloshko, and E. A. Medved. Statistical estimation of parameters for binary conditionally nonlinear autoregressive time series. Mathematical Methods of Statistics, 27(2):103–118, April 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718020023.

Proia:2018:TCR

[610] F. Proïa and M. Soltane. A test of correlation in the random coefficients of an autoregressive process. Mathematical Methods of Statistics, 27(2):119–144, April 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718020035.

Wiklund:2018:DIR

[611] T. Wiklund. The deficiency introduced by resampling. Mathematical Methods of Statistics, 27(2):145–161, April 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718020047.

Belitser:2018:LIP

[612] E. Belitser and N. Nurushev. Local inference by penalization method for biclustering model. Mathematical Methods of Statistics, 27(3):163–183, July 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718030018.

Bibi:2018:APQ

[613] A. Bibi. Asymptotic properties of QML estimation of multivariate periodic CCC–GARCH models. Mathematical Methods of Statistics, 27 (3):184–204, July 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S106653071803002X.

Golubev:2018:MAE

[614] Yu. Golubev. A minimax approach to errors-in-variables linear models. Mathematical Methods of Statistics, 27(3):205–225, July 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530718030031.

Politis:2018:TER

[615] D. N. Politis, V. A. Vasiliev, and S. E. Vorobeychikov. Truncated estimation of ratio statistics with application to heavy tail distributions. Mathematical Methods of Statistics, 27(3):226–243, July 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530718030043.

Levit:2018:OCI

[616] B. Levit. On optimal cardinal interpolation. Mathematical Methods of Statistics, 27(4):245–267, October 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530718040014.

Caron:2018:ADL

[617] E. Caron and S. Dede. Asymptotic distribution of least squares estimators for linear models with dependent errors: Regular designs. Mathematical Methods of Statistics, 27(4):268–293, October 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530718040026.

Boldin:2018:EDF

[618] M. V. Boldin and M. N. Petriev.

On the empirical distribution function of residuals in autoregression with outliers and Pearson’s chi-square type tests. Mathematical Methods of Statistics, 27(4):294–311, October 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530718040038.

Berckmoes:2018:ABC

[619] B. Berckmoes and G. Molenberghs. On the asymptotic behavior of the contaminated sample mean. Mathematical Methods of Statistics, 27(4):312–323, October 2018. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071804004X.

Hamura:2019:BPD

[620] Y. Hamura and T. Kubokawa. Bayesian predictive distribution for a negative binomial model. Mathematical Methods of Statistics, 28(1):1–17, January 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719010010.

Havet:2019:DER

[621] A. Havet, M. Lerasle, and É. Moulines.

Density estimation for RWRE. Mathematical Methods of Statistics, 28 (1):18–38, January 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530719010022.

Khardani:2019:SPM

[622] S. Khardani. A semi-parametric mode regression with censored data. Mathematical Methods of Statistics, 28(1):39–56, January 2019. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530719010034.

Boldin:2019:PPT

[623] M. V. Boldin. On the power of Pearson’s test under local alternatives in autoregression with outliers. Mathematical Methods of Statistics, 28(1):57–65, January 2019. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530719010046.

Joutard:2019:LDA

[624] C. Joutard. A large deviation approximation for multivariate density functions. Mathematical Methods of Statistics, 28(1):66–73, January 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S1066530719010058.

Klebanov:2019:OOH

[625] L. Klebanov and I. Volchenkova. Outliers and the ostensibly heavy tails. Mathematical Methods of Statistics, 28(1):74–81, January 2019. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

                                        http://link.springer.com/article/
                                        10.3103/S106653071901006X.

Balan:2019:ATL

[626] R. M. Balan and D. Jankovic. Asymptotic theory for longitudinal data with missing responses adjusted by inverse probability weights. Mathematical Methods of Statistics, 28(2):83–103, April 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719020017.

Kutta:2019:EPR

[627] T. Kutta, N. Bissantz, J. Chown, and H. Dette. The empirical process of residuals from an inverse regression. Mathematical Methods of Statistics, 28(2):104–126, April 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719020029.

LMoudden:2019:PDE

[628] A. L’Moudden and È. Marchand. On predictive density estimation under αdivergence loss. Mathematical Methods of Statistics, 28(2):127–143, April 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719020030.

Boldin:2019:APT

[629] M. V. Boldin. On the asymptotic power of tests of fit under local alternatives in autoregression. Mathematical Methods of Statistics, 28(2):144–154, April 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719020042.

Golubev:2019:MHT

[630] G. Golubev and M. Safarian. A multiple hypothesis testing approach to detection changes in distribution. Mathematical Methods of Statistics, 28(2):155–167, April 2019. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719020054.

Bouzebda:2019:CLT

[631] S. Bouzebda and B. Nemouchi. Central limit theorems for conditional empirical and conditional U-processes of stationary mixing sequences. Mathematical Methods of Statistics, 28 (3):169–207, July 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL http:

//link.springer.com/article/10.
3103/S1066530719030013.

Rimal:2019:DDS

[632] R. Rimal and M. Pensky. Density deconvolution with small Berkson errors. Mathematical Methods of Statistics, 28(3):208–227, July 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719030025.

Autin:2019:MPV

[633] F. Autin, M. Clausel, J.-M. Freyermuth, and C. Marteau. Maxiset point of view for signal detection in inverse problems. Mathematical Methods of Statistics, 28(3):228–242, July 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719030037.

Tsai:2019:AIT

[634] Ming-Tien Tsai. Admissibility of invariant tests for means with covariates. Mathematical Methods of Statistics, 28(4):243–261, October 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S106653071904001X.

Eberl:2019:SOV

[635] A. Eberl and B. Klar. On the skewness order of van Zwet and Oja. Mathematical Methods of Statistics, 28(4):262–278, October 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719040021.

Overgaard:2019:SOP

[636] M. Overgaard. State occupation probabilities in non–Markov models. Mathematical Methods of Statistics, 28(4):279–290, October 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719040033.

Khardani:2019:REP

[637] S. Khardani. Relative error prediction for twice censored data. Mathematical Methods of Statistics, 28 (4):291–306, October 2019. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719040045.

Pinelis:2019:AOT

[638] I. Pinelis. An asymptotically optimal transform of Pearson’s correlation statistic. Mathematical Methods of Statistics, 28(4):307–318, October 2019. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL

http://link.springer.com/article/
10.3103/S1066530719040057.

Comte:2020:OAE

[639] F. Comte, C. Duval, and O. Sacko. Optimal adaptive estimation on R or R^{+} of the derivatives of a density. Mathematical Methods of Statistics, 29(1):1–31, January 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720010020.

Halconruy:2020:KSN

[640] H. Halconruy and N. Marie. Kernel selection in nonparametric regression. Mathematical Methods of Statistics, 29(1):32–56, January 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720010044.

Golubev:2020:MLB

[641] Yu. Golubev and C. Pouet. Multilevel Bayes and MAP monotonicity testing. Mathematical Methods of Statistics, 29(1):57–75, January 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720010032.

Schreuder:2020:BES

[642] N. Schreuder. Bounding the expectation of the supremum of empirical processes indexed by Hölder classes. Mathematical Methods of Statistics, 29(1):76–86, January 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720010056.

Chzhen:2020:ORN

[643] Evgenii Chzhen. Optimal rates for nonparametric F -score binary classification via post-processing. Mathematical Methods of Statistics, 29 (2):87–105, April 2020. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720020027.

Schluttenhofer:2020:AMT

[644] Sandra Schluttenhofer and Jan Johannes. Adaptive minimax testing for circular convolution. Mathematical Methods of Statistics, 29 (2):106–133, April 2020. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720020039.

Mansourvar:2020:TDD

[645] Zahra Mansourvar and Majid Asadi. On a time dependent divergence measure between two residual lifetime distributions. Mathematical Methods of Statistics, 29(3):135–148, July 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530720030023.

Tavangar:2020:SMO

[646] Mahdi Tavangar and Ismihan Bayramoglu. On some models of ordered random variables and characterizations of distributions. Mathematical Methods of Statistics, 29(3):149–158, July 2020. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530720030035.

Melnikov:2020:ASF

[647] A. K. Melnikov. Applying the solution for the first multiplicity of types equation to calculate exact approximations of the probability distributions of statistical values. Mathematical Methods of Statistics, 29(4):160–171, October 2020. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530720040031.

Dupuy:2020:CGR

[648] Jean-François Dupuy. Censored gamma regression with uncertain censoring status. Mathematical Methods of Statistics, 29(4):172–196, October 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S106653072004002X.

Rukhin:2020:SAR

[649] A. L. Rukhin. Selecting an augmented random effects model. Mathematical Methods of Statistics, 29 (4):197–212, October 2020. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530720040043.

Davydov:2021:IRN

[650] Youri Davydov and Francesca Greselin. Inferential results for a new inequality curve. Mathematical Methods of Statistics, 30(1–2):1–15, January 2021. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530721010026.

Maillard:2021:LDK

[651] Odalric-Ambrym Maillard. Local Dvoretzky–Kiefer–Wolfowitz confidence bands. Mathematical Methods of Statistics, 30(1–2):16–46, January 2021. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530721010038.

Rychlik:2021:BEB

[652] Tomasz Rychlik. Bounds on the expectations of L-statistics from iid symmetric populations in various scale units. Mathematical Methods of Statistics, 30(3–4):48–63, July 2021. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530721030030.

Al-Labadi:2021:NBN

[653] Luai Al-Labadi, Forough Fazeli Asl, and Zahra Saberi. A necessary Bayesian nonparametric test for assessing multivariate normality. Mathematical Methods of Statistics, 30(3–4):64–81, July 2021. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530721030029.

Heidari:2022:DOD

[654] Maliheh Heidari, Md Abu Manju, and Edwin R. van den Heuvel. Doptimal designs for the Mitscherlich non-linear regression function. Mathematical Methods of Statistics, 31 (1):1–17, March 2022. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530722010033.

Caro-Lopera:2022:MVD

[655] Francisco J. Caro-Lopera, Graciela González Farías, and N. Balakrishnan. Matrix variate distribution theory under elliptical models — V: The non-central Wishart and inverted Wishart distributions. Mathematical Methods of Statistics, 31(1):18–42, March 2022. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530722010021.

Rychlik:2022:BES

[656] Tomasz Rychlik. Bounds on the expectations of L-statistics from iid symmetric populations in various scale units. Mathematical Methods of Statistics, 31(1):48–63, March 2022. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530721030030.

Al-Labadi:2022:NBN

[657] Luai Al-Labadi, Forough Fazeli Asl, and Zahra Saberi. A necessary Bayesian nonparametric test for assessing multivariate normality. Mathematical Methods of Statistics, 31(1):64–81, March 2022. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530721030029.

Buono:2022:VPL

[658] Francesco Buono, Maria Longobardi, and Franco Pellerey. Varentropy of past lifetimes. Mathematical Methods of Statistics, 31(2):57–73, June 2022. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S106653072202003X.

Angelini:2022:JIC

[659] Pierpaolo Angelini and Fabrizio Maturo.

Jensen’s inequality connected with a double random good. Mathematical Methods of Statistics, 31 (2):74–90, June 2022. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722020028.

Ali:2022:SIZ

[660] Essoham Ali, Mamadou Lamine Diop, and Aliou Diop. Statistical inference in a zero-inflated Bell regression model. Mathematical Methods of Statistics, 31(3):91–104, September 2022. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722030012.

ElMoumen:2022:RMA

[661] AbdelKader El Moumen, Salim Benslimane, and Samir Rahmani. Robbins– Monro algorithm with ψ-mixing random errors. Mathematical Methods of Statistics, 31(3):105–119, September 2022. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722030024.

Zamani:2022:IGF

[662] Zohreh Zamani, Omid Kharazmi, and Narayanaswamy Balakrishnan. Information generating function of record values. Mathematical Methods of Statistics, 31(3):120–133, September 2022. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722030036.

Tchazino:2022:TQE

[663] Tchamiè Tchazino, Sophie Dabo-Niang, and Aliou Diop. Tail and quantile estimation for real-valued β-mixing spatial data. Mathematical Methods of Statistics, 31(4):135–164, December 2022. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722040044.

Pinelis:2022:WIC

[664] Iosif Pinelis. What intraclass covariance structures can symmetric Bernoulli random variables have? Mathematical Methods of Statistics, 31 (4):165–169, December 2022. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722040020.

Sun:2022:TMD

[665] Ning Sun, Chen Yang, and Ricardas Zitikis. Tail maximal dependence in bivariate models: Estimation and applications. Mathematical Methods of Statistics, 31(4):170–196, December 2022. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530722040032.

Buzun:2023:GAP

[666] Nazar Buzun. Gaussian approximation for penalized Wasserstein barycenters. Mathematical Methods of Statistics, 32(1):1–26, March 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723010039.

Bouzebda:2023:RSU

[667] Salim Bouzebda and Nourelhouda Taachouche. Rates of the strong uniform consistency for the kernel-type regression function estimators with general kernels on manifolds. Mathematical Methods of Statistics, 32 (1):27–80, March 2023. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723010027.

Zohra:2023:RBD

[668] Megraoui Fatima Zohra and Belaloui Soheir. Reliability bounds of dependent circular consecutive k-out-of-n: G systems. Mathematical Methods of Statistics, 32(1):81–87, March 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723010052.

Sathar:2023:QBG

[669] E. I. Abdul Sathar and Veena L.

Vijayan. Quantile based geometric vitality function of order statistics. Mathematical Methods of Statistics, 32(1):88–101, March 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723010040.

Dijoux:2023:DDC

[670] Yann Dijoux. Distributions derived from the continuous iteration of the hyperbolic sine function. Mathematical Methods of Statistics, 32 (2):103–121, June 2023. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530723020023.

Papadatos:2023:DAT

[671] Nickos Papadatos. A discrete analogue of Terrell’s characterization of rectangular distributions. Mathematical Methods of Statistics, 32(2):122–132, June 2023. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530723020035.

Barry:2023:IET

[672] Mamadou Aliou Barry, El Hadji Deme, Aba Diop, and Solym M. Manou-Abi. Improved estimators of tail index and extreme quantiles under dependence serials. Mathematical Methods of Statistics, 32(2):133–153, June 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530723020011.

Han:2023:MDT

[673] Xiangyu Han and Chuancun Yin. Multivariate doubly truncated moments for a class of multivariate locationscale mixture of elliptical distributions. Mathematical Methods of Statistics, 32 (3):155–175, September 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723030043.

Chacko:2023:IGF

[674] Manoj Chacko and Annie Grace. Information generating function of krecord values and its applications. Mathematical Methods of Statistics, 32 (3):176–196, September 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S106653072303002X.

Shekarabi:2023:NSS

[675] F. Hosseini Shekarabi and M. Khodabin.

Numerical solution of stochastic mixed Volterra–Fredholm integral equations driven by space–time Brownian motion via two–dimensional triangular functions. Mathematical Methods of Statistics, 32(3):197–208, September 2023. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723030055.

Efromovich:2023:SLB

[676] Sam Efromovich. Sharp lower bound for regression with measurement errors and its implication for ill-posedness of functional regression. Mathematical Methods of Statistics, 32(3): 209–221, September 2023. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723030031.

Ahangari:2023:VDL

[677] Maryam Ahangari, Mousa Golalizadeh, and Zahra Rezaei Ghahroodi. Validation data-located modification for the multilevel analysis of miscategorized nominal response with covariates subject to measurement error. Mathematical Methods of Statistics, 32 (4):223–240, December 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723040026.

Amani:2023:SIM

[678] Kouakou Mathias Amani, Ouagnina Hili, and Konan Jean Geoffroy Kouakou. Statistical inference in marginalized zero-inflated Poisson regression models with missing data in covariates. Mathematical Methods of Statistics, 32 (4):241–259, December 2023. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530723040038.

Balakrishnan:2023:ESS

[679] Narayanaswamy Balakrishnan, Jan Rychtár, and Dewey Taylor. Estimating sample skewness from sample data summaries and associated evaluation of normality. Mathematical Methods of Statistics, 32(4):260–273, December 2023. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S106653072304004X.

Pham:2024:CSG

[680] Huong T. T. Pham and Hoa Pham. Controlling separation in generating samples for logistic regression models. Mathematical Methods of Statistics, 33(1):1–11, March 2024. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530724700017.

Haddouche:2024:TEP

[681] Anis M. Haddouche and Dominique Fourdrinier. Truncated estimators for a precision matrix. Mathematical Methods of Statistics, 33(1):12–25, March 2024. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530724700029.

Berrahou:2024:FUB

[682] Nour-Eddine Berrahou, Salim Bouzebda, and Lahcen Douge. Functional uniformin-bandwidth moderate deviation principle for the local empirical processes involving functional data. Mathematical Methods of Statistics, 33 (1):26–69, March 2024. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530724700030.

Bedbur:2024:CEL

[683] Stefan Bedbur, Anton Imm, and Udo Kamps. Characterizing existence and location of the ML estimate in the Conway–Maxwell–Poisson model. Mathematical Methods of Statistics, 33(1):70–78, March 2024. CO-DEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

//link.springer.com/article/10.
3103/S1066530724700042.

Chen:2024:AMA

[684] Aleksandr Chen, Nadezhda Gribkova, and Ricardas Zitikis. Assessing monotonicity: An approach based on transformed order statistics. Mathematical Methods of Statistics, 33 (1):79–94, March 2024. CODEN ???? ISSN 1066-5307 (print), 1934-8045 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        3103/S1066530724700054.